### v1

# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

Abstract
In this paper, we investigate the finite-sample statistical rates of distributional temporal difference (TD) learning with linear function approximation. The aim of distributional TD learning is to estimate the return distribution of a discounted Markov decision process for a given policy $\pi$. Prior works on statistical analysis of distributional TD learning mainly focus on the tabular case. In contrast, we first consider the linear function approximation setting and derive sharp finite-sample rates. Our theoretical results demonstrate that the sample complexity of linear distributional TD learning matches that of the classic linear TD learning. This implies that, with linear function approximation, learning the full distribution of the return using streaming data is no more difficult than learning its expectation (i.e. the value function). To derive tight sample complexity bounds, we conduct a fine-grained analysis of the linear-categorical Bellman equation, and employ the exponential stability arguments for products of random matrices. Our findings provide new insights into the statistical efficiency of distributional reinforcement learning algorithms.
1 Introduction
Distributional policy evaluation (Morimura et al., 2010; Bellemare et al., 2017, 2023), which aims to estimate the return distribution of a policy in a Markov decision process (MDP), is crucial for many uncertainty-aware or risk-sensitive tasks (Lim and Malik, 2022; Kastner et al., 2023). Unlike classic policy evaluation that only focuses on expected returns (i.e. value functions), distributional policy evaluation captures uncertainty and risk by considering the full distributional information. To solve distributional policy evaluation problem, the seminal work of Bellemare et al. (2017) proposed distributional temporal difference (TD) learning, which can be viewed as an extension of classic TD learning (Sutton, 1988).
While classic TD learning was extensively studied (Bertsekas and Tsitsiklis, 1995; Tsitsiklis and Van Roy, 1996; Bhandari et al., 2018; Dalal et al., 2018; Patil et al., 2023; Li et al., 2024a, b; Chen et al., 2024; Samsonov et al., 2024a, b; Wu et al., 2024), the theoretical understanding of distributional TD learning remains relatively underdeveloped. Recent works Rowland et al. (2018); Böck and Heitzinger (2022); Zhang et al. (2023); Rowland et al. (2024a, b); Peng et al. (2024) have analyzed distributional TD learning (or its model-based variants) in the tabular setting. Especially, Rowland et al. (2024b) and Peng et al. (2024) demonstrated that in the tabular setting, statistically speaking, learning the return distribution (in terms of the $1$-Wasserstein distance [^1] Solving distributional policy evaluation $\varepsilon$-accurately in the $1$-Wasserstein distance sense is harder than solving classic policy evaluation $\varepsilon$-accurately, as the absolute difference of value functions is always bounded by the $1$-Wasserstein distance between return distributions. See (Appendix B Peng et al., 2024) for more details.) is as easy as learning its expectation. However, in practical scenarios, where the state space is often extremely large or continuous, function approximation (Bellemare et al., 2017; Dabney et al., 2018b, a; Yang et al., 2019; Freirich et al., 2019; Yue et al., 2020; Nguyen-Tang et al., 2021; Zhou et al., 2021; Luo et al., 2022; Wenliang et al., 2024; Sun et al., 2024) becomes indispensable. This raises a new open question: When function approximation is introduced, does learning the return distribution remain as statistically efficient as learning its expectation?
To answer this question, we consider the simplest form of function approximation, i.e.  linear function approximation, and investigate the finite-sample performance of linear distributional TD learning. In distributional TD, we need to represent the infinite-dimensional return distributions with some finite-dimensional parametrizations to make the algorithm tractable. Prior works (Bellemare et al., 2019; Lyle et al., 2019; Bellemare et al., 2023) have proposed various linear distributional TD learning algorithms under different parameterizations, namely, categorical and quantile parametrizations. In this paper, we consider the categorical parametrization and propose an improved version of categorical linear TD learning algorithm (Linear-CTD). We then analyze the non-asymptotic convergence rate of Linear-CTD. Our analysis reveals that, with Polyak-Ruppert tail averaging (Ruppert, 1988; Polyak and Juditsky, 1992) and constant step size, the sample complexity of Linear-CTD matches that of the classic linear TD learning (Linear-TD) (Li et al., 2024b; Samsonov et al., 2024b). This result confirms that learning the return distribution is statistically no more difficult than learning its expectation when linear function approximation is introduced.
Notations.
In the following parts of the paper, $(x)_{+}:=\max\left\{{x,0}\right\}$ for any $x\in{\mathbb{R}}$. $\lesssim$ (resp. $\gtrsim$) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously. The asymptotic notation $f(\cdot)=\widetilde{O}\left({g(\cdot)}\right)$ (resp. $\widetilde{\Omega}\left({g(\cdot)}\right)$) means that $f(\cdot)$ is order-wise no larger (resp. smaller) than $g(\cdot)$, ignoring logarithmic factors of polynomials of $(1-\gamma)^{-1},\lambda_{\min}^{-1},\alpha^{-1}$, $\varepsilon^{-1}$, $\delta^{-1}$, $K$, $\left\|{{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$, $\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$. We will explain the meaning of the notations once we first encounter them. In addition,re the sample complexity more carefully, sometimes we will not ignore the polynomials of specific logarithmic terms in $\widetilde{O}\left({\cdot}\right)$ (resp. $\widetilde{\Omega}\left({\cdot}\right)$), we will emphasize this point when necessary.
We denote by $\otimes$ the Kronecker product (see Appendix A for properties of Kronecker product), $\bm{1}_{K}\in{\mathbb{R}}^{K}$ the all-ones vector, $\bm{0}_{K}\in{\mathbb{R}}^{K}$ the all-zeros vector, ${\bm{I}}_{K}\in{\mathbb{R}}^{K\times K}$ the identity matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of any vector ${\bm{u}}$, $\left\|{{\bm{B}}}\right\|$ the spectral norm of any matrix ${\bm{B}}$. Let $\left\|{{\bm{u}}}\right\|_{B}:=\sqrt{{\bm{u}}^{\intercal}{\bm{B}}{\bm{u}}}$ for any vector ${\bm{u}}$ and positive semi-definite (PSD) matrix ${\bm{B}}$. $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}-\bm{B}_{1}$ is PSD for any symmetric matrices $\bm{B}_{1},\bm{B}_{2}$. And $\prod_{k=1}^{t}\bm{B}_{k}$ is defined as $\bm{B}_{t}\bm{B}_{t-1}\cdots\bm{B}_{1}$ for any matrices $\left\{{\bm{B}_{k}}\right\}_{k=1}^{t}$. For any matrix ${\bm{B}}=({\bm{b}}(1),\cdots,{\bm{b}}(n))\in{\mathbb{R}}^{m\times n}$, we define its vectorization as $\mathsf{vec}({\bm{B}})=\left({{\bm{b}}(1)^{\intercal},\cdots,{\bm{b}}(n)^{\intercal}}\right)^{\intercal}\in{\mathbb{R}}^{mn}$. Suppose $A$ is some set. We denote by $\Delta(A)$ the set of all probability distributions over $A$. And we denote $\Delta\left({\left[{0,(1-\gamma)^{-1}}\right]}\right)$ by ${\mathscr{P}}$.
1.1 Contributions
Our contribution includes two folds.
Algorithmically, we propose an improved version of linear categorical TD learning algorithm (Linear-CTD). Instead of using stochastic semi-gradient descent (SSGD) to update the parameter as in (Bellemare et al., 2019; Lyle et al., 2019; Bellemare et al., 2023), we directly translate the linear-categorical projected Bellman equation into a linear system and apply linear stochastic approximation (LSA) to solve it. The resulting Linear-CTD can be viewed as a preconditioned version of the vanilla linear categorical TD learning algorithm proposed in (Section 9.6 Bellemare et al., 2023). By introducing a preconditioner, our Linear-CTD achieve a finite-sample rate independent of the number of supports $K$ in the categorical parameterization, a result the vanilla form cannot attain. We also note that our proposed Linear-CTD is similar to the algorithm in (Lyle et al., 2019), with the distinction that our approach ensures the total mass of the estimated return distributions always be $1$.
Theoretically, we establish the first non-asymptotic guarantees for distributional TD learning with linear function approximation. Specifically, we show that, in the generative model setting, with Polyak-Ruppert tail averaging and instance-dependent constant step size,
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}}\left({\frac{1}{\varepsilon^{2}}+\frac{1}{\lambda_{\min}}}\right)}\right)$
[…]
online interactions with the environment are sufficient to guarantee Linear-CTD to yield an $\varepsilon$-accurate estimator with high probability, when the error is measured by the $\mu_{\pi}$-weighted $1$-Wasserstein distance. We also provide sample complexity bounds when we take the instance-independent (i.e. not dependent on unknown quantity) step size, and in the Markovian setting. Our sample complexity bounds match those of classic Linear-TD with constant step size (Li et al., 2024b; Samsonov et al., 2024b), confirming the equal statistical tractability of distributional and classic value-based policy evaluation. To establish these theoretical results, we analyze the linear-categorical Bellman equation in details, and apply the exponential stability argument proposed in (Samsonov et al., 2024b). Our analysis of the linear-categorical Bellman equation lays the foundation for subsequent algorithmic and theoretical advances in distributional reinforcement learning with function approximation.
1.2 Related Work
Distributional Reinforcement Learning.
Distributional TD learning was first proposed in (Bellemare et al., 2017). Following the distributional perspective in (Bellemare et al., 2017), Qu et al. (2019) proposed a distributional version of the gradient TD learning algorithm, Tang et al. (2022) proposed a distributional version of multi-step TD learning, Tang et al. (2024) proposed a distributional version of off-policy Q($\lambda$) and TD($\lambda$) algorithms, and Wu et al. (2023) proposed a distributional version of fitted Q evaluation to solve the distributional offline policy evaluation problem. Wiltzer et al. (2024b) proposed an approach for evaluating the return distributions for all policies simultaneously when the reward is deterministic or in the finite-horizon setting. Wiltzer et al. (2024a) studied distributional policy evaluation in the multivariate reward setting and proposed corresponding TD learning algorithms. Beyond the tabular setting, Bellemare et al. (2019); Lyle et al. (2019); Bellemare et al. (2023) proposed various distributional TD learning algorithms with linear function approximation under different parametrizations.
A series of recent studies have focused on the theoretical properties of distributional TD learning. Rowland et al. (2018); Böck and Heitzinger (2022); Zhang et al. (2023); Rowland et al. (2024a, b); Peng et al. (2024) analyzed the asymptotic and non-asymptotic convergence of distributional TD learning (or its model-based variants) in the tabular setting. Among these works, Rowland et al. (2024b); Peng et al. (2024) established that in the tabular setting, learning the full return distribution is statistically as easy as learning its expectation, in the model-based and model-free setting respectively. And Bellemare et al. (2019) provided an asymptotic convergence result for categorical TD learning with linear function approximation.
Beyond the problem of distributional policy evaluation, Rowland et al. (2023); Wang et al. (2023, 2024) showed that, theoretically, classic value-based reinforcement learning could benefit from distributional reinforcement learning. Bäuerle and Ott (2011); Chow and Ghavamzadeh (2014); Marthe et al. (2023); Noorani et al. (2023); Pires et al. (2025) considered optimizing statistical functionals of the return, and proposed algorithms to solve this harder problem.
Stochastic Approximation.
Our Linear-CTD falls into the category of LSA. The classic TD learning, as one of the most classic LSA problems, has been extensively studied (Bertsekas and Tsitsiklis, 1995; Tsitsiklis and Van Roy, 1996; Bhandari et al., 2018; Dalal et al., 2018; Patil et al., 2023; Duan and Wainwright, 2023; Li et al., 2024a, b; Samsonov et al., 2024a; Wu et al., 2024). Among these works, Li et al. (2024b); Samsonov et al. (2024b) provided the tightest bounds for Linear-TD with constant step sizes, which is also considered in our paper. While Wu et al. (2024) established the tightest bounds for Linear-TD with polynomial-decaying step sizes.
For general stochastic approximation problems, extensive works including (Lakshminarayanan and Szepesvari, 2018; Srikant and Ying, 2019; Mou et al., 2020, 2022; Huo et al., 2023; Li et al., 2023b; Durmus et al., 2024; Samsonov et al., 2024b; Chen et al., 2024) have provided solid theoretical understandings.
1.3 Organization
The remainder of this paper is organized as follows. In Section 2, we recap Linear-TD, categorical parametrization of return distributions and categorical TD learning. In Section 3, we introduce the linear-categorical parametrization, and use the linear-categorical projected Bellman equation to derive the Linear-CTD algorithm. In Section 4, we employ the exponential stability arguments to analyze the statistical efficiency of Linear-CTD in both the generative model setting and the Markovian setting. In Section 6, we conclude our work. The details of our proof are given in the appendices.
2 Backgrounds
2.1 Policy Evaluation and TD learning
A discounted MDP is defined by a 5-tuple $M=\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}_{R},P,\gamma\rangle$. We assume the state space ${\mathcal{S}}$ and action space ${\mathcal{A}}$ are both Polish spaces, namely a complete separable metric space. ${{\mathcal{P}}_{R}}$ is the distribution of rewards, i.e. ${\mathcal{P}}_{R}(\cdot|s,a)\in\Delta\left({[0,1]}\right)$ for any state-action pair $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ (we assume all rewards is bounded). ${P}$ is the transition dynamics, i.e. $P(\cdot|s,a)\in\Delta\left({{\mathcal{S}}}\right)$ for any $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$. And $\gamma\in(0,1)$ is the discount factor.
Given a policy $\pi\colon{\mathcal{S}}\to\Delta\left({{\mathcal{A}}}\right)$ and an initial state $s_{0}=s\in{\mathcal{S}}$, a random trajectory $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$ can be sampled: $a_{t}\mid s_{t}\sim\pi(\cdot\mid s_{t})$, $r_{t}\mid(s_{t},a_{t})\sim{\mathcal{P}}_{R}({\cdot}\mid s_{t},a_{t})$, ${s_{t+1}}\mid{(s_{t},a_{t})}\sim P({\cdot}\mid{s_{t},a_{t}})$ for any $t\in{\mathbb{N}}$. We assume the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ has a unique stationary distribution $\mu_{\pi}\in\Delta({\mathcal{S}})$. We define the return of the trajectory by $G^{\pi}(s):=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The value function $V^{\pi}(s)$ is the expectation of $G^{\pi}(s)$, and ${\bm{V}}^{\pi}:=\left({V^{\pi}(s)}\right)_{s\in{\mathcal{S}}}$. It is known that ${\bm{V}}^{\pi}$ satisfies the Bellman equation:
$V^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),r\sim{\mathcal{P}}_{R}({\cdot}\mid s,a),s^{\prime}\sim P(\cdot\mid s,a)}\left[{r+\gamma V^{\pi}(s^{\prime})}\right],\quad\forall s\in{\mathcal{S}},$ (1)
or in a compact form ${\bm{V}}^{\pi}={\bm{T}}^{\pi}{\bm{V}}^{\pi}$, where ${\bm{T}}^{\pi}\colon{\mathbb{R}}^{\mathcal{S}}\to{\mathbb{R}}^{\mathcal{S}}$ is called the Bellman operator.
In the task of policy evaluation, we aim to find ${\bm{V}}^{\pi}$ for some given policy $\pi$. Since ${\bm{V}}^{\pi}$ is the unique solution of the Bellman equation, the problem is reduced to solving the Bellman equation. However, in practical applications ${\bm{T}}^{\pi}$ is usually unknown and the agent only has access to the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$. In this circumstance, we may solve the Bellman equation through LSA. Specifically, at the $t$-th time-step, the updating scheme is
${V}_{t}(s_{t})=V_{t-1}(s_{t})-\alpha\left({V_{t-1}(s_{t})-r_{t}-\gamma V_{t-1}(s_{t+1})}\right).$
We expect ${\bm{V}_{t}}$ will converge to ${\bm{V}}^{\pi}$ as $t$ tends to infinity. This algorithm is known as the TD learning.
2.2 Linear Function Approximation and Linear TD Learning
In this section, we introduce linear function approximation and briefly review Linear-TD. To be concrete, we assume there is a $d$-dimensional feature for each state $s\in{\mathcal{S}}$, which is given by the feature map ${\bm{\phi}}\colon{\mathcal{S}}\to{\mathbb{R}}^{d}$. We consider the linear function approximation of value functions:
${\mathscr{V}}_{{\bm{\phi}}}:=\left\{{\bm{V}}_{{\bm{\psi}}}=\left({V_{{\bm{\psi}}}(s)}\right)_{s\in{\mathcal{S}}}\colon V_{{\bm{\psi}}}(s)={\bm{\psi}}^{\intercal}{\bm{\phi}}(s),{\bm{\psi}}\in{\mathbb{R}}^{d}\right\}\subset{\mathbb{R}}^{\mathcal{S}},$
$\mu_{\pi}$-weighted norm $\left\|{{\bm{V}}}\right\|_{\mu_{\pi}}:=({\mathbb{E}}_{s\sim\mu_{\pi}}[V(s)^{2}])^{1/2}$, and linear projection operator ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\colon{\mathbb{R}}^{{\mathcal{S}}}\to{\mathscr{V}}_{{\bm{\phi}}}$:
${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}{\bm{V}}:=\operatorname*{argmin}\nolimits_{{\bm{V}}_{{\bm{\psi}}}\in{\mathscr{V}}_{{\bm{\phi}}}}\left\|{{\bm{V}}-{\bm{V}}_{{\bm{\psi}}}}\right\|_{\mu_{\pi}},\quad\forall{\bm{V}}\in{\mathbb{R}}^{\mathcal{S}}.$
One can check that the linear projected Bellman operator ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ is a $\gamma$-contraction in the Polish space $({\mathscr{V}}_{{\bm{\phi}}},\left\|{\cdot}\right\|_{\mu_{\pi}})$. Hence, ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ admits a unique fixed point ${\bm{V}}_{{\bm{\psi}}^{\star}}$, which satisfies $\|{\bm{V}}^{\pi}-{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}\leq(1-\gamma^{2})^{-1/2}\|{\bm{V}}^{\pi}-{\bm{\Pi}}_{{\bm{\phi}}}^{\pi}{\bm{V}}^{\pi}\|_{\mu_{\pi}}$ (Theorem 9.8 Bellemare et al., 2023). In Appendix B.1, we show that ${\bm{\psi}}^{\star}\in{\mathbb{R}}^{d}$ is the unique solution to the linear system for ${\bm{\psi}}\in{\mathbb{R}}^{d}$:
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right].$ (2)
In the subscript of the expectation, we abbreviate $s\sim\mu_{\pi}(\cdot),a\sim\pi(\cdot|s),r\sim{\mathcal{P}}_{R}(\cdot|s,a),s^{\prime}\sim P(\cdot|s,a)$ as $s,a,r,s^{\prime}$. For brevity, we will continue to use such abbreviations in this paper when there is no ambiguity. We can use LSA to solve the linear projected Bellman equation (Eqn. (2)). As a result, at the $t$-th time-step, the updating scheme of Linear-TD is
${\bm{\psi}}_{t}={\bm{\psi}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left[{\left({{\bm{\phi}}(s_{t})-\gamma{\bm{\phi}}(s_{t+1})}\right)^{\intercal}{\bm{\psi}}_{t-1}-r_{t}}\right].$ (3)
See Appendix B.2 for convergence results for Linear-TD.
2.3 Distributional Policy Evaluation and Distributional TD Learning
In certain applications, we are not only interested in finding the expectation of the random return $G^{\pi}(s)$ but want to find the whole distribution of $G^{\pi}(s)$. This task is called distributional policy evaluation. We use $\eta^{\pi}(s)\in{\mathscr{P}}$ to denote the distribution of $G^{\pi}(s)$ and let ${\bm{\eta}}^{\pi}:=(\eta^{\pi}(s))_{s\in{\mathcal{S}}}\in{\mathscr{P}}^{\mathcal{S}}$. Then ${\bm{\eta}}^{\pi}$ satisfies the distributional Bellman equation:
$\eta^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),r\sim{\mathcal{P}}_{R}({\cdot}\mid s,a),s^{\prime}\sim P(\cdot\mid s,a)}[\left({b_{r,\gamma}}\right)_{\#}\eta^{\pi}(s^{\prime})],\quad\forall s\in{\mathcal{S}}.$ (4)
Here $b_{r,\gamma}\colon{\mathbb{R}}\to{\mathbb{R}}$ is the affine function defined by $b_{r,\gamma}(x)=r+\gamma x$, and $f_{\#}\nu$ is the push forward measure of $\nu$ through any function $f\colon{\mathbb{R}}\to{\mathbb{R}}$, so that $f_{\#}\nu(A)=\nu(f^{-1}(A))$ for any Borel set $A$, where $f^{-1}(A):=\left\{{x\colon f(x)\in A}\right\}$. The distributional Bellman equation can also be written as ${\bm{\eta}}^{\pi}={\bm{\bm{\mathcal{T}}}}^{\pi}{\bm{\eta}}^{\pi}$. The operator ${{\bm{\mathcal{T}}}}^{\pi}\colon{\mathscr{P}}^{\mathcal{S}}\to{\mathscr{P}}^{\mathcal{S}}$ is called the distributional Bellman operator and ${\bm{\eta}}^{\pi}$ is the associated fixed point.
In analogy to TD learning, we can solve the distributional Bellman equation by LSA and get the distributional TD learning algorithm given the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$:
$\eta_{t}(s_{t})=\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})].$
We comment the algorithm above is not computationally feasible as we need to manipulate an infinite-dimensional object at each iteration.
2.4 Categorical Parametrization of Return Distributions and Categorical TD Learning
In order to deal with the infinite-dimensional return distributions in a computational tractable manner, we consider the finite-dimensional categorical parametrization as in (Bellemare et al., 2017; Rowland et al., 2018; Bellemare et al., 2019; Rowland et al., 2024b; Peng et al., 2024). To incorporate function approximation in the distribution parametrization which cannot guarantee non-negative outputs, we will work with the signed measure space ${\mathscr{P}}^{\mathsf{sign}}$ with total mass $1$ instead of the standard probability space ${\mathscr{P}}\subset{\mathscr{P}}^{\mathsf{sign}}$:
${\mathscr{P}}^{\mathsf{sign}}:=\left\{\nu\colon\left|{\nu}\right|({\mathbb{R}})<\infty,\nu({\mathbb{R}})=1,\operatorname{supp}(\nu)\subseteq\left[0,(1-\gamma)^{-1}\right]\right\}.$
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, we define its cumulative distribution function (CDF) as $F_{\nu}(x):=\nu[0,x]$. We can naturally define the $L^{2}$ and $L^{1}$ distance between CDFs as Cramér distance $\ell_{2}$ and $1$-Wasserstein distance $W_{1}$ in ${\mathscr{P}}^{\mathsf{sign}}$ respectively. The distributional Bellman operator (see Eqn. (4)) can also be extended to the product space $({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$ without modifying its definition.
The space of all categorical parametrized signed measures with total mass $1$ is defined as
${\mathscr{P}}^{\mathsf{sign}}_{K}:=\left\{\nu_{\bm{p}}=\sum_{k=0}^{K}p_{k}\delta_{x_{k}}\colon{\bm{p}}=\left({p_{0},\cdots,p_{K-1}}\right)^{\intercal}\in{\mathbb{R}}^{K},p_{K}=1-\sum_{k=0}^{K-1}p_{k}\right\},$
which is an affine subspace of ${\mathscr{P}}^{\mathsf{sign}}$, where $K\in{\mathbb{N}}$, $\left\{{x_{k}=k\iota_{K}}\right\}_{k=0}^{K}$ are equally-spaced points of the support, and $\iota_{K}=\left[{K(1-\gamma)}\right]^{-1}$ is the gap between two adjacent points. We say ${\bm{p}}$ is the probability mass function (PMF) representation of $\nu_{\bm{p}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$. In fact, this representation is an isometry (see Proposition B.1). We define the categorical projection operator ${\bm{\Pi}}_{K}\colon{\mathscr{P}}^{\mathsf{sign}}\to{\mathscr{P}}^{\mathsf{sign}}_{K}$ as
${\bm{\Pi}}_{K}\nu:=\operatorname*{argmin}\nolimits_{\nu_{\bm{p}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}}\ell_{2}\left({\nu,\nu_{\bm{p}}}\right),\quad\forall\nu\in{\mathscr{P}}^{\mathsf{sign}}.$
Following (Proposition 5.14 Bellemare et al., 2023), one can show that $\bm{\Pi}_{K}\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ is uniquely identified with a vector ${\bm{p}}_{\nu}=\left({p_{k}(\nu)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, where
$p_{k}(\nu)={\mathbb{E}}_{X\sim\nu}[(1-\left|{(X-x_{k})/{\iota_{K}}}\right|)_{+}]:=\int_{\left[{0,(1-\gamma)^{-1}}\right]}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx).$ (5)
$\bm{\Pi}_{K}$ is in fact an orthogonal projection (see Proposition B.2), and thus is non-expansive.
For any ${\bm{\eta}}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{\mathcal{S}}$, $s\in{\mathcal{S}}$, we define ${\bm{p}}_{\bm{\eta}}(s):={\bm{p}}_{\eta(s)}$ and lift ${\bm{\Pi}}_{K}$ to the product space by defining $\left[{\bm{\Pi}_{K}{\bm{\eta}}}\right](s):=\bm{\Pi}_{K}\eta(s)$. One can check that the categorical Bellman operator ${\bm{\Pi}}_{K}{\bm{\mathcal{T}}}^{\pi}$ (see Proposition B.3 for the characterization of $\bm{\Pi}_{K}\bm{\mathcal{T}}^{\pi}$) is a $\sqrt{\gamma}$-contraction in the Polish space $(({\mathscr{P}}_{K}^{\mathsf{sign}})^{\mathcal{S}},\ell_{2,\mu_{\pi}})$, where $\ell_{2,\mu_{\pi}}(\bm{\eta}_{1},\bm{\eta}_{2}):=({\mathbb{E}}_{s\sim\mu_{\pi}}[\ell_{2}^{2}(\eta_{1}(s),\eta_{2}(s))])^{1/2}$ is the $\mu_{\pi}$-weighted Cramér distance between $\bm{\eta}_{1},\bm{\eta}_{2}\in({\mathscr{P}}^{\mathsf{sign}})^{\mathcal{S}}$. Hence, the categorical projected Bellman equation $\bm{\eta}={\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}\bm{\eta}$ admits a unique solution $\bm{\eta}^{\pi,K}$, which satisfies $\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})\leq(1-\gamma)^{-1/2}\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},{\bm{\Pi}}_{K}\bm{\eta}^{\pi})$ (Proposition 3 Rowland et al., 2018). Applying LSA to solving the equation yields categorical TD learning, and the iteration rule is given by
$\eta_{t}(s_{t})=\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-{\bm{\Pi}}_{K}\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})].$
3 Linear-Categorical TD Learning
3.1 Linear-Categorical Parametrization of Return Distributions
Now, we combine linear function approximation with categorical parametrization, and introduce the space of linear-categorical parametrized signed measures with total mass $1$: [^3] Compared with the parametrization in (Section 9.6 Bellemare et al., 2023), our model is identifiable because we remove redundant parameters in their model, which can be linearly represented by other parameters.
${\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}:=\left\{\bm{\eta}_{{\bm{w}}}=\left({\eta_{{\bm{w}}}(s)}\right)_{s\in{\mathcal{S}}}\colon\eta_{{\bm{w}}}(s)=\sum_{k=0}^{K}p_{k}(s;{\bm{w}})\delta_{x_{k}},{\bm{w}}\in{\mathbb{R}}^{dK}\right\},$
which is an affine subspace of $({\mathscr{P}}^{\mathsf{sign}}_{K})^{{\mathcal{S}}}$, where ${\bm{w}}=({\bm{w}}(0)^{\intercal},\cdots,{\bm{w}}(K-1)^{\intercal})^{\intercal}$ and
$p_{k}(s;{\bm{w}})=\begin{cases}{\bm{\phi}}(s)^{\intercal}{\bm{w}}(k)+\frac{1}{K+1},&\text{ for}\ k\in\left\{{0,1,\cdots K-1}\right\};\\ -{\bm{\phi}}(s)^{\intercal}\sum_{j=0}^{K-1}{\bm{w}}(j)+\frac{1}{K+1},&\text{ for}\ k=K.\end{cases}$ (6)
Again, this representation is an isometry (see Proposition C.1). In many cases, it is more convenient to work with the matrix version of the parameter ${\bm{W}}:=\left({{\bm{w}}(0),\cdots,{\bm{w}}(K-1)}\right)\in{\mathbb{R}}^{d\times K}$ instead of ${\bm{w}}=\mathsf{vec}\left({{\bm{W}}}\right)$. We abbreviate ${\bm{p}}_{\bm{\eta}_{\bm{w}}}$ as ${\bm{p}}_{{\bm{w}}}$, then by Lemma A.2, for any $s\in{\mathcal{S}}$, it holds that
${\bm{p}}_{{\bm{w}}}(s)={\bm{W}}^{\intercal}{\bm{\phi}}(s)+(K+1)^{-1}\bm{1}_{K}.$ (7)
We define the linear-categorical projection operator ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\colon({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}\to{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$ as follow:
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}:=\operatorname*{argmin}\nolimits_{\bm{\eta}_{{\bm{w}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}}\ell_{2,\mu_{\pi}}\left({\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right),\quad\forall\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}.$
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$ is in fact an orthogonal projection (see Proposition C.2), and thus is non-expansive. The following proposition characterizes ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$, whose proof can be found in Appendix C.2.
Proposition 3.1.
For any $\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$, ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}$ is uniquely give by $\bm{\eta}_{\tilde{\bm{w}}}$, where
$\tilde{{\bm{w}}}=\mathsf{vec}(\tilde{\bm{W}}),\quad\tilde{\bm{W}}={\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}[{\bm{\phi}}(s)\left({{\bm{p}}_{\bm{\eta}}(s)-(K+1)^{-1}\bm{1}_{K}}\right)^{\intercal}].$
3.2 Linear-Categorical Projected Bellman Equation
Since ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in the Polish space $({\mathscr{P}}_{{\bm{\phi}},K}^{\mathsf{sign}},\ell_{2,\mu_{\pi}})$, we have the following theorem, whose proof can be found in Appendix C.3.
Theorem 3.1.
The linear-categorical projected Bellman equation $\bm{\eta}_{{\bm{w}}}={\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{w}}}$ admits a unique solution $\bm{\eta}_{{\bm{w}}^{\star}}$, where the matrix version of the parameter ${\bm{W}}^{\star}\in{\mathbb{R}}^{d\times K}$ is the unique solution to the linear system for ${\bm{W}}\in{\mathbb{R}}^{d\times K}$
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}-{\mathbb{E}}_{s,s^{\prime},r}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]=\frac{1}{K+1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$ (8)
where for any $r\in[0,1]$ and $j,k\in\left\{{0,1\cdots,K}\right\}$,
$g_{j,k}(r):=\left({1-\left|{(r+\gamma x_{j}-x_{k})/{\iota_{K}}}\right|}\right)_{+},\quad{\bm{g}}_{j}(r):=\left({g_{j,k}(r)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K},$
${\bm{G}}(r):=\begin{bmatrix}{\bm{g}}_{0}(r),&\cdots,&{\bm{g}}_{K-1}(r)\end{bmatrix}\in{\mathbb{R}}^{K\times K},\quad\tilde{{\bm{G}}}(r):={\bm{G}}(r)-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r)\in{\mathbb{R}}^{K\times K}.$
In analogy to approximation results of $\|{\bm{V}}^{\pi}-{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}$ and $\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})$, the following lemma answers how close $\bm{\eta}_{{\bm{w}}^{\star}}$ is to $\bm{\eta}^{\pi}$, whose proof can be found in Appendix C.5.
Proposition 3.2 (Approximation Error of $\bm{\eta}_{{\bm{w}}^{\star}}$).
Consider the $\mu_{\pi}$-weighted $1$-Wasserstein error $W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right):=({\mathbb{E}}_{s\sim\mu_{\pi}}[W_{1}^{2}({\eta}^{\pi}(s),{\eta}_{{\bm{w}}^{\star}}(s))])^{1/2}$, it holds that
$W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right)\leq$ $K^{-1}(1-\gamma)^{-3}+(1-\gamma)^{-2}\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right),$
where the first error term $K^{-1}(1-\gamma)^{-3}$ is due to the categorical parametrization, and the second error term $(1-\gamma)^{-2}\ell_{2,\mu_{\pi}}^{2}({\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi})$ is due to the additional linear function approximation.
3.3 Update Scheme of Linear-Categorical TD Learning
As before, we solve Eqn. (8) by LSA and get Linear-CTD given the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$:
${\bm{W}}_{t}{=}{\bm{W}}_{t{-}1}{-}\alpha{\bm{\phi}}(s_{t})\!\!\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t{-}1}{-}{\bm{\phi}}(s_{t{+}1})^{\intercal}{\bm{W}}_{t{-}1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K{+}1}\!\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K}}\right)^{\intercal}}\right],$ (9)
for any $t\geq 1$, where $\alpha$ is the constant step size. In this paper, we also consider the Polyak-Ruppert tail averaging $\bar{{\bm{w}}}_{T}:=\left({T/2+1}\right)^{-1}\sum_{t=T/2}^{T}{\bm{w}}_{t}$ as in the analysis of Linear-TD in (Samsonov et al., 2024b). Standard theory of LSA (Mou et al., 2020) tells us, under some conditions, if we take an appropriate constant step size, $\bar{{\bm{w}}}_{T}$ will converges to the solution ${\bm{w}}^{\star}$ with rate $T^{-1/2}$ as $T\to\infty$.
Remark 1 (Comparison with Existing Works).
Our Linear-CTD can be regarded as a preconditioned version of SSGD with the PMF representation considered in (Bellemare et al., 2019) and (Section 9.6 Bellemare et al., 2023). The preconditioning technique (Chen, 2005; Li, 2017) is a commonly used methodology to accelerate solving optimization problems by reducing condition number. We precondition the vanilla SSGD with the PMF representation by removing the matrix ${\bm{C}}^{\intercal}{\bm{C}}$ on the right of Eqn. (23), whose condition number scales with $K^{2}$ (Lemma E.2). By introducing the preconditioner $\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1}$, our Linear-CTD (Eqn. (9)) can achieve a convergence rate independent of $K$, which the vanilla form cannot achieve (see Appendix D.6). It is worth noting that our Linear-CTD (Eqn. (9)) is equivalent to the stochastic semi-gradient descent (SSGD) with CDF representation, which was also considered in (Lyle et al., 2019). See Appendix C.6 for the CDF representation, and Appendix C.7 for a self-contained derivation of SSGD with different representations. We further comment that our Linear-CTD can guarantee that the total mass of return distributions always be $1$, while the algorithms proposed in (Lyle et al., 2019) cannot.
4 Non-Asymptotic Statistical Analysis
In the task of distributional policy evaluation, the quality of estimator $\bm{\eta}_{\bar{\bm{w}}_{T}}$ can be measured by $\mu_{\pi}$-weighted $1$-Wasserstein error $W_{1,\mu_{\pi}}(\bm{\eta}_{\bar{\bm{w}}_{T}},\bm{\eta}^{\pi})$. By triangle inequality, the error can be decomposed as approximation error and estimation error: $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{\bar{\bm{w}}_{T}})\leq W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}})+W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{w}}^{\star}},\bm{\eta}_{\bar{\bm{w}}_{T}})$. Proposition 3.2 provides an upper bound for the approximation error $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}})$, so it suffices to control the estimation error ${\mathcal{L}}(\bar{\bm{w}}_{T}):=W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{w}}^{\star}},\bm{\eta}_{\bar{\bm{w}}_{T}})$.
In the following, we give non-asymptotic convergence rates of ${\mathcal{L}}(\bar{\bm{w}}_{T})$. We start from the generative model setting, i.e. in the $t$-th iteration, we collect samples $s_{t}\sim\mu_{\pi}(\cdot),a_{t}\sim\pi(\cdot|s_{t}),r_{t}\sim{\mathcal{P}}_{R}(\cdot|s_{t},a_{t}),s_{t}^{\prime}\sim P(\cdot|s_{t},a_{t})$ from the generative model (we need to replace $s_{t+1}$ with $s_{t}^{\prime}$ in Linear-CTD Eqn. (9)). We give $L^{p}$ and high-probability convergence results in this setting. Then, we move back to the Markovian setting (i.e.  using the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$ introduced in Section 2.1), and give high-probability convergence results.
To facilitate the comparison of our results with those of Linear-TD, in the following, we denote by ${\bm{\theta}}_{0}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}_{0}$ (resp. ${\bm{\theta}}^{\star}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}^{\star}$) the initial (resp. optimal) parameters under CDF representation, where ${\bm{C}}$ is defined in Eqn. (16). See Appendix C.6 for details of CDF representation.
4.1 $L^{2}$ Convergence
We first provide non-asymptotic convergence rates of ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]$.
Theorem 4.1 ($L^{2}$ Convergence).
Suppose $K\geq(1-\gamma)^{-1}$, $T\geq 2$, $\alpha\in(0,(1-\sqrt{\gamma})/76)$ and ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$ is the initialization, then it holds that
${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]\lesssim$ $\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\sqrt{T}(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)+\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{T\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{T\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|.$
To prove Theorem 4.1, we conduct a fine-grained analysis of the linear-categorical Bellman equation and apply the exponential stability argument (Theorem 1 Samsonov et al., 2024b), which is outlined in Section 5. And the $L^{p}$ ($p>2$) convergence results can be found in Theorem D.1. Theorem 4.1 implies that learning the distribution of the return is as easy as learning its expectation (value function) with linear function approximation. Rowland et al. (2024b); Peng et al. (2024) discovered this phenomenon in the tabular setting, and we extend it to the function approximation setting.
Remark 2.
The only difference between our Theorem 4.1 and the $L^{2}$ convergence rate of classic Linear-TD (Theorem 3 Samsonov et al., 2024b) lies in replacing $\left\|{\bm{\psi}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$ (resp. $\left\|{\bm{\psi}_{0}{-}\bm{\psi}^{\star}}\right\|$) with $K^{-1/2}(1{-}\gamma)^{-1}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$ (resp. $K^{-1/2}(1{-}\gamma)^{-1}\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|$). We claim that the two pairs should be of the same order respectively. Note that $\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|$ are of the order $O(\sqrt{K})$ (also dependent on ${\bm{\phi}}$, we omit it for brevity) if ${\eta}_{{\bm{w}}^{\star}}(s),{\eta}_{{\bm{w}}_{0}}(s)$ are valid probability distributions for all $s\in{\mathcal{S}}$. This is because in this case, the vector of CDF ${\bm{F}}_{{\bm{\theta}}}(s)=({\bm{\theta}}(k)^{\intercal}{\bm{\phi}}(s){+}({k{+}1})/({K{+}1}))_{k=0}^{K{-}1}\in\left[{0,1}\right]^{K}$ for ${\bm{\theta}}\in\left\{{{\bm{\theta}}^{\star},{\bm{\theta}}_{0}}\right\}$. While, in classic Linear-TD (Section 2.2), a proper estimate $\bm{\psi}$ should satisfy $V_{\bm{\psi}}(s)={\bm{\psi}}^{\intercal}{\bm{\phi}}(s)\in\left[{0,(1{-}\gamma)^{-1}}\right]$. It is thus reasonable to consider the two pairs as being of the same order, respectively. Similar arguments also holds in other convergence results presented in this paper. Therefore, in this sense, our results match those of classic Linear-TD.
One can translate Theorem 4.1 into a sample complexity bound.
Corollary 4.1.
Under the same conditions as Theorem 4.1, for any $\varepsilon>0$, suppose
$T\gtrsim$ $\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\left({1+\frac{\alpha}{(1-\gamma)\lambda_{\min}}}\right)+\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{1}{\alpha(1-\gamma)\lambda_{\min}}\left({\log\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|}{\varepsilon\sqrt{K}(1{-}\gamma)}+\log\left({\frac{1}{T\sqrt{\alpha}(1{-}\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)}\right)}\right),$
then it holds that ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]\leq\varepsilon$.
Instance-Independent Step Size.
If we take the largest possible instance-independent step size, i.e. $\alpha\simeq(1-\gamma)$, and consider $\varepsilon\in(0,1)$, we obtain the sample complexity bound
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}^{2}}}\right).$ (10)
Optimal Instance-Dependent Step Size.
If we take the optimal instance-dependent step size $\alpha\simeq(1-\gamma)\lambda_{\min}$ which involves the unknown $\lambda_{\min}$, we obtain a better sample complexity bound
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}}\right),$ (11)
when we consider small enough $\varepsilon=\tilde{O}\left({\sqrt{\lambda_{\min}}}\right)$, or equivalently, large enough total update steps
$T=\widetilde{\Omega}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}^{2}}}\right).$ (12)
These theoretical results match the recent results for classic Linear-TD with constant step size (Patil et al., 2023; Li et al., 2024b; Samsonov et al., 2024b). It is possible to break the sample size barrier (Eqn. (12)) as in classic Linear-TD by applying the variance-reduction techniques (Li et al., 2023a) , we leave it as a future work.
4.2 Convergence with High Probability and Markovian Samples
By applying the $L^{p}$ convergence result (Theorem D.1) with $p=2\log(1/\delta)$ and Markov’s inequality, we obtain the high-probability convergence result.
Theorem 4.2 (High-Probability Convergence).
For any $\varepsilon>0$ and $\delta\in(0,1)$, suppose $K\geq(1{-}\gamma)^{-1}$, $\alpha\in(0,(1{-}\sqrt{\gamma}){/}[38\log(T{/}\delta^{2})])$, ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$ is the initialization, and total update steps $T=$
$\widetilde{O}\Bigg{(}$ $\frac{\frac{1}{K(1{-}\gamma)^{2}}\!\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}{\otimes}{\bm{\Sigma}}_{{\bm{\phi}}}}\!\!\!{+}1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\!\left({1{+}\frac{\alpha\log\frac{1}{\delta}}{(1{-}\gamma)\lambda_{\min}}}\right)\!\log\!\frac{1}{\delta}{+}\frac{\frac{1}{\sqrt{K}(1{-}\gamma)}\!\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}{\otimes}{\bm{\Sigma}}_{{\bm{\phi}}}}\!\!\!{+}1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}\!\log\!\frac{1}{\delta}{+}\frac{\log\!\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|}{\varepsilon\sqrt{K}(1{-}\gamma)}}{\alpha(1{-}\gamma)\lambda_{\min}}\Bigg{)},$
then with probability at least $1-\delta$, it holds that ${\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)\leq\varepsilon$. Here, the $\widetilde{O}\left({\cdot}\right)$ does not hide polynomials of $\log(1/\delta)$ (but hides logarithm terms of $\log(1/\delta)$).
Again, we will obtain concrete sample complexity bounds as in Eqn. (10) or Eqn. (11) if we use different step sizes, we omit these for brevity. Compared with the theoretical results for classic Linear-TD, our results match (Theorem 4 Samsonov et al., 2024b) that also considers the constant step size, but has a worse dependence on $\log\left({1/\delta}\right)$ than (Theorem 3.1 Wu et al., 2024) which considers the polynomial-decaying step size $\alpha_{t}=\alpha_{0}t^{-\beta}$ with $\beta\in(1/2,1)$ instead.
Remark 3 (Markovian Setting).
Using the same argument as in the proof of (Theorem 6 Samsonov et al., 2024b), one can immediately derive a high-probability sample complexity bound in the Markovian setting. Compared with the bound in the generative model setting (Theorem 4.2), the bound in the Markovian setting will have an additional dependency on $t_{\operatorname{mix}}\log(T/\delta)$, where $t_{\operatorname{mix}}$ is the mixing time of the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ in ${\mathcal{S}}$. We omit this result for brevity.
5 Proof Outlines
In this section, we outline the proofs of our main results (Theorem 4.1). We first state the theoretical properties of the linear-categorical Bellman equation and the exponential stability of Linear-CTD. Finally, we highlight some key steps in proving these results.
5.1 Vectorization of Linear-CTD
In our analysis, it will be more convenient to work with the vectorization version and the CDF representation introduced in Appendix C.6. In Appendix C.7, we show that Linear-CTD (Eqn. (9)) is equivalent to SSGD with the CDF representation (Eqn. (24)). We consider the vectorization of Eqn. (24), and we denote ${\bm{\theta}}_{t}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}_{t}$ (other CDF parameters are defined in the same way):
${\bm{\theta}}_{t}{=}{\bm{\theta}}_{t-1}{-}\alpha\left({{\bm{A}}_{t}{\bm{\theta}}_{t-1}{-}{\bm{b}}_{t}}\right),\quad{\bm{A}}_{t}{=}\left[{{\bm{I}}_{K}{\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s_{t})^{\intercal}}\right)}\right]{-}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1}}\right){\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s^{\prime}_{t})^{\intercal}}\right)}\right],$
${\bm{b}}_{t}=\frac{1}{K+1}\left[{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s_{t}).$ (13)
We denote by $\bar{{\bm{A}}},\bar{{\bm{b}}}$ the expectation of ${\bm{A}}_{t},{\bm{b}}_{t}$ respectively. Utilizing the exponential stability arguments, we can derive an upper bound for $\left\|{\bar{{\bm{A}}}\left({\bar{{\bm{\theta}}}_{T}{-}{\bm{\theta}}^{\star}}\right)}\right\|$. We need to further translate it to an upper bound for ${\mathcal{L}}(\bar{{\bm{w}}}_{T})$, which is done in the following lemma, whose proof can be found in Appendix D.1.
Lemma 5.1.
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$, it holds that ${\mathcal{L}}({\bm{w}})\leq 2{K^{-1/2}(1{-}\gamma)^{-2}\lambda_{\min}^{-1/2}}\left\|{\bar{{\bm{A}}}\left({{{\bm{\theta}}}{-}{\bm{\theta}}^{\star}}\right)}\right\|$.
5.2 Exponential Stability Analysis
First, we introduce some notations. Let ${\bm{e}}_{t}:{=}{{\bm{A}}}_{t}{\bm{\theta}}^{\star}{-}{{\bm{b}}}_{t}$, we denote by $C_{A}$ (resp. $C_{e}$) the almost sure upper bound for $\max\left\{{\left\|{{\bm{A}}_{t}}\right\|,\left\|{{\bm{A}}_{t}{-}\bar{{\bm{A}}}}\right\|}\right\}$ (resp. $\left\|{{\bm{e}}_{t}}\right\|$), and ${\bm{\Sigma}}_{e}:={\mathbb{E}}\left[{{\bm{e}}_{t}{\bm{e}}_{t}^{\intercal}}\right]$ the covariance matrix of ${\bm{e}}_{t}$. The following lemma provides useful upper bounds, whose proof can be found in Appendix D.2.
Lemma 5.2.
Suppose $K\geq(1-\gamma)^{-1}$, then it holds that
$C_{A}\leq 4,\quad C_{e}\leq 4(\left\|{{\bm{\theta}}^{\star}}\right\|+\sqrt{K}\left({1-\gamma}\right)),\quad\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)\leq 18(\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}+K(1-\gamma)^{2}).$
Let ${\bm{\Gamma}}_{m:n}^{(\alpha)}:=\prod_{i=m}^{n}\left({{\bm{I}}-\alpha{\bm{A}}_{i}}\right)$ for any $\alpha>0$ and $m,n\in{\mathbb{N}}$, $m\leq n$. The exponential stability of Linear-CTD is summarized in the following lemma, whose proof can be found in Appendix D.3.
Lemma 5.3.
For any $p\geq 2$, let $a=(1-\sqrt{\gamma})\lambda_{\min}/2$ and $\alpha_{p,\infty}=(1-\sqrt{\gamma})/(38p)$ ($\alpha_{p,\infty}p\leq 1/2$). Then for any $\alpha\in\left({0,\alpha_{p,\infty}}\right)$, ${\bm{u}}\in{\mathbb{R}}^{dK}$ and $n\in{\mathbb{N}}$, it holds that ${\mathbb{E}}^{1/p}[\|{\bm{\Gamma}}_{1:n}^{(\alpha)}{\bm{u}}\|^{p}]\leq(1-\alpha a)^{n}\left\|{{\bm{u}}}\right\|$.
See Appendix D.6 for the counterparts of Lemma 5.1, Lemma 5.2 and Lemma 5.3 for vanilla SSGD with the PMF representation, which have additional dependency on $K$.
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with (Theorem 1 Samsonov et al., 2024b), one can immediately obtain Theorem 4.1, the remaining details can be found in Appendix D.4.
5.3 Key Steps in the Proofs
In this section, we highlight some key steps in proving above theoretical results.
Bounding Spectral Norm of Expectation of Kronecker Product.
In proving that the ${\mathcal{L}}({\bm{w}})$ can be upper-bounded by $\left\|{\bar{{\bm{A}}}\left({{{\bm{\theta}}}-{\bm{\theta}}^{\star}}\right)}\right\|$ (Lemma 5.1), as well as in verifying the exponential stability condition (Lemma 5.3), one of the most critical steps is to show
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma},\quad\forall r\in[0,1].$ (14)
By Lemma E.3, we have $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ for any $r\in[0,1]$. In addition, one can check that $\|{\mathbb{E}}_{s,s^{\prime}}[{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}]\|{\leq}1$. One may speculate that the property $\left\|{{\bm{B}}_{1}{\otimes}{\bm{B}}_{2}}\right\|=\left\|{{\bm{B}}_{1}}\right\|\left\|{{\bm{B}}_{2}}\right\|$ (Lemma A.3) is enough to get the desired conclusion. However, the two matrices in the Kronecker product are not independent, preventing us from using this simple property to derive the conclusion. On the other hand, since we merely have the upper bound ${\mathbb{E}}_{s,s^{\prime}}[\|{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}\|]{\leq}d$, simply moving the expectation in Eqn. (14) outside the norm will lead to a loose $d\sqrt{\gamma}$ bound . To resolve this problem, we leverage the fact that the second matrix is rank-$1$ and prove the following result. The proof can be found in the derivation following Eqn. (26).
Lemma 5.4.
For any random matrix ${\bm{Y}}$ and random vectors ${\bm{x}}$, ${\bm{z}}$, suppose $\left\|{{\bm{Y}}}\right\|\leq C_{Y}$ almost surely, ${\mathbb{E}}\left[{{\bm{x}}{\bm{x}}^{\intercal}}\right]\preccurlyeq C_{x}{\bm{I}}_{d_{1}}$ and ${\mathbb{E}}\left[{{\bm{z}}{\bm{z}}^{\intercal}}\right]\preccurlyeq C_{z}{\bm{I}}_{d_{2}}$ for some constants $C_{Y},C_{x},C_{z}>0$, then it holds that
$\left\|{{\mathbb{E}}\left[{{\bm{Y}}\otimes\left({{\bm{x}}{\bm{z}}^{\intercal}}\right)}\right]}\right\|\leq C_{Y}\sqrt{C_{x}C_{z}}.$
Remark 4.
The matrix ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ also appeared in (Proposition B.2 Rowland et al., 2024b) as the matrix representation of the categorical projected Bellman operator ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ of an one-state MDP. They showed $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ using the fact that ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in $\left({{\mathscr{P}},\ell_{2}}\right)$. Our Lemma E.3 provides a new analysis by directly analyzing the matrix. See Appendix F for more details.
Bounding Norm of ${\bm{b}}_{t}$.
In proving Lemma 5.2, the most involved step is to upper-bound $\left\|{{\bm{b}}_{t}}\right\|$ (Eqn. (13)). To this end, we need to upper-bound the following term which appears in Eqn. (27):
$(K+1)^{-1}\|\sum_{j=0}^{K}{\bm{C}}{\bm{g}}_{j}(r)-{\bm{C}}\bm{1}_{K}\|,\quad\forall r\in[0,1],$ (15)
when $K\geq(1-\gamma)^{-1}$. Same as the matrix ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ (see Remark 4), Term (15) is also related to the matrix representation of the categorical projected Bellman operator. Specifically, let $\nu=(K+1)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ be the discrete uniform distribution, Term (15) equals
$\left\|{{\bm{C}}\left({{\bm{p}}_{(b_{r,\gamma})_{\#}(\nu)}-{\bm{p}}_{\nu}}\right)}\right\|=\iota_{K}^{-1/2}\ell_{2}\left({{\bm{\Pi}}_{K}(b_{r,\gamma})_{\#}(\nu),\nu}\right)\leq\iota_{K}^{-1/2}\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)\leq 3\sqrt{K}(1-\gamma),$
where we used the orthogonal decomposition (Proposition B.2) and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ (Lemma E.4). The full proof can be found in the derivation following Eqn. (27).
6 Conclusions
In this paper, we have bridged a critical theoretical gap in distributional reinforcement learning by establishing the non-asymptotic sample complexity of distributional TD learning with linear function approximation. Specifically, we have proposed Linear-CTD, derived via solving the linear-categorical projected Bellman equation with linear stochastic approximation. By carefully analyzing the Bellman equation and using the exponential stability arguments, we have shown tight sample complexity bounds for the proposed algorithm in both generative model setting and Markovian setting. Our finite-sample rates match the state-of-the-art sample complexity bounds for the classic TD learning. These theoretical findings demonstrate that learning the full return distribution under linear function approximation can be statistically as easy as classic TD learning for value function estimation. Finally, we noted that it is possible to improve the convergence rates by applying variance-reduction techniques or using polynomial-decaying step sizes, which we leave as a future work.
Appendix A Kronecker Product
In this section, we will introduce some properties of Kronecker product used in our paper. See [Zhang and Ding, 2013] for a detailed treatment of Kronecker product.
For any matrices ${\bm{A}}\in{\mathbb{R}}^{m\times n}$ and ${\bm{B}}\in{\mathbb{R}}^{p\times q}$, the Kronecker product ${\bm{A}}\otimes{\bm{B}}$ is an matrix in ${\mathbb{R}}^{mp\times nq}$, defined as
${\bm{A}}\otimes{\bm{B}}=\begin{bmatrix}a_{11}{\bm{B}}&a_{12}{\bm{B}}&\cdots&a_{1n}{\bm{B}}\\ a_{21}{\bm{B}}&a_{22}{\bm{B}}&\cdots&a_{2n}{\bm{B}}\\ \vdots&\vdots&\ddots&\vdots\\ a_{m1}{\bm{B}}&a_{m2}{\bm{B}}&\cdots&a_{mn}{\bm{B}}\end{bmatrix}.$
Lemma A.1.
The Kronecker product is bilinear and associative. Furthermore, for any matrices ${\bm{B}}_{1},{\bm{B}}_{2},{\bm{B}}_{3},{\bm{B}}_{4}$ such that ${\bm{B}}_{1}{\bm{B}}_{3}$, ${\bm{B}}_{2}{\bm{B}}_{4}$ can be defined, it holds that $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)=\left({{\bm{B}}_{1}{\bm{B}}_{3}}\right)\otimes\left({{\bm{B}}_{2}{\bm{B}}_{4}}\right)$ (mixed-product property).
Proof.
See [Basic properties and Theorem 3 Zhang and Ding, 2013]. ∎
Lemma A.2.
For any matrices ${\bm{B}}_{1},{\bm{B}}_{2},{\bm{B}}_{3}$ such that ${\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}$ can be defined, it holds that $\mathsf{vec}\left({{\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}}\right)=\left({{\bm{B}}_{3}^{\intercal}\otimes{\bm{B}}_{1}}\right)\mathsf{vec}\left({{\bm{B}}_{2}}\right)$.
Proof.
See [Lemma 4.3.1 Horn and Johnson, 1994]. ∎
Lemma A.3.
For any matrices ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$, it holds that $\left\|{{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right\|=\left\|{{\bm{B}}_{1}}\right\|\left\|{{\bm{B}}_{2}}\right\|$, $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{\intercal}={\bm{B}}_{1}^{\intercal}\otimes{\bm{B}}_{2}^{\intercal}$. Furthermore, if ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$ are invertible/orthogonal/diagonal/symmetric/normal, ${\bm{B}}_{1}\otimes{\bm{B}}_{2}$ is also invertible/orthogonal/diagonal/symmetric/normal and $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{-1}={\bm{B}}_{1}^{-1}\otimes{\bm{B}}_{2}^{-1}$.
Proof.
See [Basic properties, Theorem 5 and Theorem 7 Zhang and Ding, 2013]. ∎
Lemma A.4.
For any $K,d\in{\mathbb{N}}$ and PSD matrices ${\bm{B}}_{1},{\bm{B}}_{3}\in{\mathbb{R}}^{K\times K},{\bm{B}}_{2},{\bm{B}}_{4}\in{\mathbb{R}}^{d\times d}$ with ${\bm{B}}_{1}\preccurlyeq{\bm{B}}_{3}$ and ${\bm{B}}_{2}\preccurlyeq{\bm{B}}_{4}$, it holds that ${\bm{B}}_{1}\otimes{\bm{B}}_{2}$, ${\bm{B}}_{3}\otimes{\bm{B}}_{4}$ are also PSD matrices, furthermore, ${\bm{B}}_{1}\otimes{\bm{B}}_{2}\preccurlyeq{\bm{B}}_{3}\otimes{\bm{B}}_{4}$.
Proof.
Consider the spectral decomposition ${\bm{B}}_{i}={\bm{Q}}_{i}{\bm{D}}_{i}{\bm{Q}}_{i}^{\intercal}$, for any $i\in[4]$, by Lemma A.1 and Lemma A.3, we have
$\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)=\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)\left({{\bm{D}}_{1}\otimes{\bm{D}}_{2}}\right)\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)^{\intercal}$
and
$\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)=\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)\left({{\bm{D}}_{3}\otimes{\bm{D}}_{4}}\right)\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)^{\intercal}$
are also spectral decomposition of $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)$ and $\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)$ respectively. It is easy to see that they are PSD. Furthermore,
$\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)-\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)=$ $\left[{\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)-\left({{\bm{B}}_{3}\otimes{\bm{B}}_{2}}\right)}\right]+\left[{\left({{\bm{B}}_{3}\otimes{\bm{B}}_{2}}\right)-\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)}\right]$
$=$ $\left[{{\bm{B}}_{3}\otimes\left({{\bm{B}}_{4}-{\bm{B}}_{2}}\right)}\right]+\left[{\left({{\bm{B}}_{3}-{\bm{B}}_{1}}\right)\otimes{\bm{B}}_{2}}\right]$
[…]
Lemma A.5.
For any $K,d,d_{1},d_{2}\in{\mathbb{N}}$, vectors ${\bm{u}},{\bm{v}}\in{\mathbb{R}}^{d}$ and matrices ${\bm{B}}_{1}\in{\mathbb{R}}^{K\times d_{1}}$, ${\bm{B}}_{2}\in{\mathbb{R}}^{d_{2}\times K}$, ${\bm{B}}_{3}\in{\mathbb{R}}^{K\times K}$, it holds that
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{1}={\bm{B}}_{1}\otimes{\bm{u}},$
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}={\bm{B}}_{2}\otimes{\bm{v}}^{\intercal},$
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\intercal}}\right).$
Furthermore, for any matrix ${\bm{B}}_{4}\in{\mathbb{R}}^{d_{1}\times d_{2}}$, we have
$\left({{\bm{B}}_{1}\otimes{\bm{u}}}\right){\bm{B}}_{4}=\left({{\bm{B}}_{1}{\bm{B}}_{4}}\right)\otimes{\bm{u}}.$
Proof.
Let ${\bm{u}}=\left({u_{i}}\right)_{i=1}^{d}$ ${\bm{B}}_{1}=\left({b_{ij}}\right)_{i,j=1}^{K}$, then
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{1}=$ $\begin{bmatrix}{\bm{u}}&\bm{0}_{d}&\cdots&\bm{0}_{d}&\bm{0}_{d}\\ \bm{0}_{d}&{\bm{u}}&\cdots&\bm{0}_{d}&\bm{0}_{d}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \bm{0}_{d}&\bm{0}_{d}&\cdots&{\bm{u}}&\bm{0}_{d}\\ \bm{0}_{d}&\bm{0}_{d}&\cdots&\bm{0}_{d}&{\bm{u}}\end{bmatrix}\begin{bmatrix}b_{11}&\cdots&b_{1K}\\ \vdots&\ddots&\vdots\\ b_{K1}&\cdots&b_{KK}\end{bmatrix}$
$=$ $\begin{bmatrix}b_{11}u_{1}&\cdots&b_{1K}u_{1}\\ \vdots&\ddots&\vdots&\\ b_{11}u_{d}&\cdots&b_{1K}u_{d}\\ \vdots&\ddots&\vdots&\\ b_{K1}u_{1}&\cdots&b_{KK}u_{1}\\ \vdots&\ddots&\vdots&\\ b_{K1}u_{d}&\cdots&b_{KK}u_{d}\end{bmatrix}$
$=$ $\begin{bmatrix}b_{11}{\bm{u}}&\cdots&b_{1K}{\bm{u}}\\ \vdots&\ddots&\vdots\\ b_{K1}{\bm{u}}&\cdots&b_{KK}{\bm{u}}\end{bmatrix}$
$=$ ${\bm{B}}_{1}\otimes u.$
Hence
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}=$ $\left[{\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)\otimes{\bm{B}}^{\intercal}_{2}}\right]^{\intercal}=\left[{{\bm{B}}^{\intercal}_{2}\otimes{\bm{v}}}\right]^{\intercal}={\bm{B}}_{2}\otimes{\bm{v}}^{\intercal}.$
And in the same way,
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}=$ $\left({{\bm{B}}_{3}\otimes{\bm{u}}}\right)\otimes{\bm{v}}^{\intercal}={\bm{B}}_{3}\otimes\left({{\bm{u}}\otimes{\bm{v}}^{\intercal}}\right)={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\intercal}}\right).$
Furthermore,
$\left({{\bm{B}}_{1}\otimes{\bm{u}}}\right){\bm{B}}_{4}=\left[{\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{1}}\right]{\bm{B}}_{4}=\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right)\left({{\bm{B}}_{1}{\bm{B}}_{4}}\right)=\left({{\bm{B}}_{1}{\bm{B}}_{4}}\right)\otimes{\bm{u}}.$
Appendix B Omitted Results and Proofs in Section 2
B.1 Linear Projected Bellman Equation
It is worth noting that, ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\colon\left({{\mathbb{R}}^{\mathcal{S}},\left\|{\cdot}\right\|_{\mu_{\pi}}}\right)\to\left({{\mathscr{V}}_{\bm{\phi}},\left\|{\cdot}\right\|_{\mu_{\pi}}}\right)$ is an orthogonal projection.
We aim to derive Eqn. (2). It is easy to check that, for any ${\bm{V}}\in{\mathbb{R}}^{{\mathcal{S}}}$, ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}{\bm{V}}$ is uniquely give by ${\bm{V}}_{\tilde{\bm{\psi}}}$ where
$\tilde{\bm{\psi}}={\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)V(s)}\right].$
Hence, by the definition of Bellman operator (Eqn. (1)), ${\bm{\psi}}^{\star}$ is the unique solution to the following system of linear equations for ${\bm{\psi}}\in{\mathbb{R}}^{d}$
${\bm{\psi}}=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left[{{\bm{T}}^{\pi}{\bm{V}}_{{\bm{\psi}}}}\right](s)}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({{\mathbb{E}}\left[{r_{0}\mid s_{0}=s}\right]+\gamma{\mathbb{E}}\left[{{\bm{\phi}}(s_{1})^{\intercal}\mid s_{0}=s}\right]{\bm{\psi}}}\right)}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]{\bm{\psi}}+{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right],$
or equivalently,
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right].$
B.2 Convergence Results for Linear TD Learning
It is worthy noting that, Linear-TD is equivalent to the stochastic semi-gradient descent (SSGD) update.
In Linear-TD, our goal is to find a good estimator $\hat{{\bm{\psi}}}$ such that $\left\|{{\bm{V}}_{\hat{{\bm{\psi}}}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}=\left\|{\hat{{\bm{\psi}}}-{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}\leq\varepsilon$. Samsonov et al. [2024b] considered the Polyak-Ruppert tail averaging $\bar{{\bm{\psi}}}_{T}:=\left({T/2+1}\right)^{-1}\sum_{t=T/2}^{T}{\bm{\psi}}_{t}$, and showed that in the generative model setting with constant step size $\alpha\simeq(1-\gamma)\lambda_{\min}$,
$T=\widetilde{O}\left({\frac{\left\|{{\bm{\psi}}^{\star}}\right\|^{2}_{{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}}\left({\frac{1}{\varepsilon^{2}}+\frac{1}{\lambda_{\min}}}\right)}\right)$
is sufficient to guarantee that $\left\|{{\bm{V}}_{\bar{{\bm{\psi}}}_{T}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}\leq\varepsilon$. They also provided sample complexity bounds when taking the instance-independent (i.e. not dependent on unknown quantity) step size, and in the Markovian setting.
B.3 Categorical Parametrization is an Isometry
Proposition B.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{K},\ell_{2}}\right)$ is isometric with $\left({{\mathbb{R}}^{K},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{C}}^{\intercal}{\bm{C}}}}\right)$, in the sense that, for any $\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that $\ell_{2}^{2}(\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}})=\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\intercal}{\bm{C}}}$, where
${\bm{C}}=\begin{bmatrix}1&0&\cdots&0&0\\ 1&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&1&\cdots&1&0\\ 1&1&\cdots&1&1\end{bmatrix}\in{\mathbb{R}}^{K\times K}.$ (16)
Proof.
[…]
$=$ $\iota_{K}\sum_{k=0}^{K-1}\left({F_{\nu_{{\bm{p}}_{1}}}(x_{k})-F_{\nu_{{\bm{p}}_{2}}}(x_{k})}\right)^{2}$
$=$ $\iota_{K}\left\|{{\bm{C}}\left({{\bm{p}}_{1}-{\bm{p}}_{2}}\right)}\right\|^{2}$
$=$ $\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\intercal}{\bm{C}}}.$
B.4 Categorical Projection Operator is Orthogonal Projection
Proposition B.2.
[Bellemare et al., 2023, Lemma 9.17] For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$ and $\nu_{{\bm{p}}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that
$\ell_{2}^{2}\left({\nu,\nu_{\bm{p}}}\right)=\ell_{2}^{2}\left({\nu,\bm{\Pi}_{K}\nu}\right)+\ell_{2}^{2}\left({\bm{\Pi}_{K}\nu,\nu_{\bm{p}}}\right).$
B.5 Categorical Projected Bellman Operator
The following lemma characterizing $\bm{\Pi}_{K}\bm{\mathcal{T}}^{\pi}$ is useful for both practice and theoretical analysis.
Proposition B.3.
For any ${\bm{\eta}}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{\mathcal{S}}$ and $s\in{\mathcal{S}}$, it holds that
${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}}(s)=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{\bm{\eta}}(s_{1})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\tilde{\bm{G}}(r_{0})\left({{\bm{p}}_{\bm{\eta}}(s_{1})-\frac{1}{K+1}\bm{1}_{K}}\right)\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right].$
And in the same way, for any $r\in[0,1]$ and $s^{\prime}\in{\mathcal{S}}$, it holds that
${\bm{p}}_{\left({b_{r,\gamma}}\right)_{\#}{\eta}(s^{\prime})}$ $=\tilde{\bm{G}}(r)\left({{\bm{p}}_{\bm{\eta}}(s^{\prime})-\frac{1}{K+1}\bm{1}_{K}}\right)+\frac{1}{K+1}\sum_{j=0}^{K}{\bm{g}}_{j}(r),$
where $\tilde{{\bm{G}}}$ and ${\bm{g}}$ is defined in Theorem 3.1.
This proposition is a special case of Proposition C.3, whose proof can be found in Appendix C.4.
Appendix C Omitted Results and Proofs in Section 3
C.1 Linear-Categorical Parametrization is an Isometry
Proposition C.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with $\left({{\mathbb{R}}^{dK},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$, in the sense that, for any $\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that $\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}}\right)=\iota_{K}\left\|{{\bm{w}}_{1}-{\bm{w}}_{2}}\right\|^{2}_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$.
Proof.
By Proposition B.1,
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}}\right)=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{\bm{C}}\left({{\bm{p}}_{{\bm{w}}_{1}}(s)-{\bm{p}}_{{\bm{w}}_{2}}(s)}\right)}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right){\bm{C}}^{\intercal}{\bm{C}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}}\right)$
$=$ $\iota_{K}\left\|{{\bm{w}}_{1}-{\bm{w}}_{2}}\right\|^{2}_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}.$
C.2 Linear-Categorical Projection Operator
Proposition 3.1 is an immediate corollary of the following lemma.
Lemma C.1.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$, ${\bm{w}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, it holds that
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{W}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Furthermore, it holds that
$\nabla_{{\bm{W}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}}\right)$ $={\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)}\right]$
$=2\iota_{K}\left[{{\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}+{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Proof.
According to Proposition B.2, one has
$\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)=\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\bm{\Pi}}_{K}{\eta}(s)}\right)+\ell_{2}^{2}\left({{\bm{\Pi}}_{K}{\eta}(s),{\eta}(s)}\right).$
Hence,
$\nabla_{{\bm{w}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=\nabla_{{\bm{w}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\bm{\Pi}}_{K}{\eta}(s)}\right)$
$=\iota_{K}\nabla_{{\bm{w}}}\left\|{{\bm{C}}\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right\|^{2}$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\intercal}{\bm{C}}\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\intercal}{\bm{C}}\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left\{{\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]{\bm{w}}+\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\},$
where in the second equality, we used Proposition B.1, and in the last equality, we used Lemma A.5. We also have the following matrix representation:
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{W}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Proposition C.2.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$ and $\bm{\eta}_{{\bm{w}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right)=\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},{\bm{\Pi}}_{K}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right).$
The proof is straightforward and almost the same as that of Proposition B.2 if we utilize the affine structure.
C.3 Linear-Categorical Projected Bellman Equation
To derive the result, the following proposition characterizing ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{w}}}$ is useful, whose proof can be found in Appendix C.4.
Proposition C.3.
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, we abbreviate ${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}(s)$ as $\tilde{{\bm{p}}}_{{\bm{w}}}(s)$, then
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{w}})}\right)_{k=0}^{K-1}={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right],$
Combining this proposition with Proposition 3.1, we know that ${\bm{W}}^{\star}$ is the unique solution to the following system of linear equations for ${\bm{W}}\in{\mathbb{R}}^{d\times K}$
${\bm{W}}=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\tilde{\bm{p}}_{{\bm{w}}}(s)-\frac{1}{K+1}\bm{1}_{K}}\right)^{\intercal}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]-\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right]}\right)^{\intercal}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]+\frac{1}{K+1}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$
or equivalently,
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}-{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]=\frac{1}{K+1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$
which is the desired conclusion. The uniqueness and existence of the solution is guaranteed by the fact that the LHS is an invertible linear transformation of ${\bm{W}}$, which is justified by Eqn. (25).
C.4 Proof of Proposition C.3
Proof.
Recall the definition of the distributional Bellman operator Eqn. (4) and categorical projection operator Eqn. (5), we have
$\tilde{p}_{k}(s;{\bm{w}})=$ $p_{k}\left({\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}\right](s)}\right)$ (17)
$=$ ${\mathbb{E}}_{X\sim\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}\right](s)}\left[{\left({1-\left|{\frac{X-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ ${\mathbb{E}}\left[{{\mathbb{E}}_{G\sim\eta_{{\bm{w}}}(s_{1})}\left[{\left({1-\left|{\frac{r_{0}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{w}})\left({1-\left|{\frac{r_{0}+\gamma x_{j}-x_{k}}{\iota_{K}}}\right|}\right)_{+}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{w}})g_{j,k}(r_{0})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{g_{K,k}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\left({g_{j,k}(r_{0})-g_{K,k}(r_{0})}\right)\Big{|}s_{0}=s}\right].$
Hence,
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{w}})}\right)_{k=0}^{K-1}$
$=$ ${\mathbb{E}}\left[{\begin{bmatrix}g_{K,1}(r_{0})\\ \vdots\\ g_{K,{K-1}}(r_{0})\end{bmatrix}+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\begin{bmatrix}g_{j,1}(r_{0})-g_{K,1}(r_{0})\\ \vdots\\ g_{j,K-1}(r_{0})-g_{K,K-1}(r_{0})\end{bmatrix}\Bigg{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\left({{\bm{g}}_{j}(r_{0})-{\bm{g}}_{K}(r_{0})}\right)\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{{\bm{w}}}(s_{1})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\left[{\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\intercal}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\intercal}\Big{|}s_{0}=s}\right]{\bm{w}}$
$\quad+{\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\frac{1}{K+1}\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\bm{1}_{K}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\otimes{\bm{\phi}}(s_{1})^{\intercal}\Big{|}s_{0}=s}\right]{\bm{w}}+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right],$
or equivalently,
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right].$
C.5 Proof of Proposition 3.2
Proof.
By the basic inequality (Lemma E.1), we only need to show
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right)\leq$ $\frac{\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right)}{1-\gamma}$
$=$ $\frac{\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},{\bm{\Pi}}_{K}\bm{\eta}^{\pi}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right)}{1-\gamma}$
$\leq$ $\frac{1}{K(1-\gamma)^{2}}+\frac{\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right)}{1-\gamma},$
where we used [Proposition 9.18 and Eqn. (5.28) Bellemare et al., 2023]. ∎
C.6 Cumulative Distribution Function Representation
C.6.1 Categorical Parametrization
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, we denote
${\bm{F}}_{\nu}=\left({F_{k}\left({\nu}\right)}\right)_{k=0}^{K-1}=\left({\nu\left({[0,x_{k}]}\right)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$
as the cumulative distribution function (CDF) representation of $\nu$. One can check that ${\bm{F}}_{\nu}={\bm{C}}{\bm{p}}_{\nu}$. As a result, PMF and CDF representations are equivalent because ${\bm{C}}$ is invertible.
For any $\nu_{1},\nu_{2}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$,
$\ell_{2}^{2}(\nu_{1},\nu_{2})=$ $\iota_{K}\left\|{{\bm{F}}_{\nu_{1}}-{\bm{F}}_{\nu_{2}}}\right\|^{2},$
therefore, the affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{K},\ell_{2}}\right)$ is isometric with the Euclidean space $\left({{\mathbb{R}}^{K},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{2}}\right)$ if we consider the CDF representation.
For any ${\bm{\eta}}\in\left({{\mathscr{P}}^{\mathsf{sign}}_{K}}\right)^{\mathcal{S}}$, $s\in{\mathcal{S}}$, we abuse the notation to define ${\bm{F}}_{\bm{\eta}}(s):={\bm{F}}_{\eta(s)}$. By Eqn. (5), we can also similarly define ${\bm{F}}_{\nu},{\bm{F}}_{\bm{\eta}}(s)$ using the categorical projection operator for any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$ and $s\in{\mathcal{S}}$.
C.6.2 Linear-Categorical Parametrization
We introduce new notations for the CDF representation when we consider linear-categorical parametrization. Let ${\bm{\Theta}}:={\bm{W}}{\bm{C}}^{\intercal}=\left({\sum_{k=0}^{0}{\bm{w}}(k),\sum_{k=0}^{1}{\bm{w}}(k),\cdots,\sum_{k=0}^{K}{\bm{w}}(k)}\right)\in{\mathbb{R}}^{d\times K}$ and the vectorization of ${\bm{\Theta}}$, ${\bm{\theta}}:=\mathsf{vec}\left({{\bm{\Theta}}}\right)=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}\in{\mathbb{R}}^{dK}$. We abbreviate ${\bm{F}}_{\bm{\eta}_{\bm{w}}}$ as ${\bm{F}}_{{\bm{\theta}}}$, then by Lemma A.2 and Lemma A.5, for any $s\in{\mathcal{S}}$, it holds that
${\bm{F}}_{{\bm{\theta}}}(s)={\bm{C}}{\bm{p}}_{{\bm{w}}}(s)={\bm{\Theta}}^{\intercal}{\bm{\phi}}(s)+\frac{1}{K+1}{\bm{C}}\bm{1}_{K}=\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{\theta}}+\frac{1}{K+1}{\bm{C}}\bm{1}_{K},$ (18)
Again, PMF and CDF representations are equivalent because ${\bm{C}}$ is invertible.
For any $\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, by Proposition B.1,
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}}\right)=$ ${\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\ell_{2}^{2}\left({\eta_{{\bm{w}}_{1}}(s),\eta_{{\bm{w}}_{2}}(s)}\right)}\right]$ (19)
$=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{\bm{F}}_{{\bm{\theta}}_{1}}(s)-{\bm{F}}_{{\bm{\theta}}_{2}}(s)}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)}\right)$
$=$ $\iota_{K}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$
hence the affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with the Euclidean space $\left({{\mathbb{R}}^{Kd},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$ if we consider the CDF representation.
Following the proof of Lemma C.1 in Appendix C.2, we can also derive the gradient when we use the CDF parametrization:
$\nabla_{{\bm{\theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=\iota_{K}\nabla_{{\bm{\theta}}}\left\|{{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right\|^{2}$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)\left({{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{\theta}}+{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)$
$=2\iota_{K}\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]{\bm{\theta}}+\left[{\left({{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\},$
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right)^{\intercal}$ (20)
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Theta}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}}\right].$
C.7 Stochastic Semi-Gradient Descent with Linear Function Approximation
We use the notations in the generative model setting for simplicity. We denote by ${\bm{\mathcal{T}}}_{t}^{\pi}$ the corresponding empirical distributional Bellman operator at the $t$-th iteration, which satisfies
$\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}}}\right](s_{t})=(b_{r_{t},\gamma})_{\#}(\eta(s_{t}^{\prime})),\quad\forall\bm{\eta}\in{\mathscr{P}}^{{\mathcal{S}}}.$ (21)
C.7.1 PMF Representation
Consider the SSGD with the PMF representation
${\bm{W}}_{t}={\bm{W}}_{t-1}-\alpha\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{W}}}$ stands for taking gradient w.r.t. ${\bm{W}}_{t-1}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{w}}_{t-1}}(s_{t})$ (the second term is regarding as a constant, that’s why we call it a semi-gradient). One can check that $\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{W}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right)$.
Now, let’s compute the gradient term. By Lemma C.1, we have
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ $=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
where ${\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})={\bm{p}}_{{\bm{\Pi}}_{K}\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}(s_{t})}=\left({p_{k}\left({\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$. Now, we turn to compute ${\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})$. According to Eqn. (5),
$p_{k}\left({\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)=$ ${\mathbb{E}}_{X\sim\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\left[{\left({1-\left|{\frac{X-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ ${\mathbb{E}}_{G\sim\eta_{{\bm{w}}_{t-1}}(s_{t}^{\prime})}\left[{\left({1-\left|{\frac{r_{t}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ $\sum_{j=0}^{K}p_{j}(s_{t}^{\prime};{\bm{w}}_{t-1})g_{j,k}(r_{t})$
$=$ $g_{K,k}(r_{t})+\sum_{j=0}^{K-1}p_{j}(s_{t}^{\prime};{\bm{w}}_{t-1})\left({g_{j,k}(r_{t})-g_{K,k}(r_{t})}\right),$
which has the same form as Eqn. (17). Following the proof of Proposition C.3 in Appendix C.4, one can show that
${\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})=\tilde{{\bm{G}}}(r_{t}){\bm{W}}^{\intercal}_{t-1}{\bm{\phi}}(s_{t}^{\prime})+\frac{1}{K+1}\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}).$ (22)
Hence, the update scheme is
${\bm{W}}_{t}=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$ (23)
$=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Compared to Linear-CTD (Eqn. (9)), this form has an additional ${\bm{C}}^{\intercal}{\bm{C}}$, and the step size is $2\iota_{K}\alpha$.
C.7.2 CDF Representation
Consider the SSGD with the CDF representation
${\bm{\Theta}}_{t}={\bm{\Theta}}_{t-1}-\alpha\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{\Theta}}}$ stands for taking gradient w.r.t. ${\bm{\Theta}}_{t-1}={\bm{W}}_{t-1}{\bm{C}}^{\intercal}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{w}}_{t-1}}(s_{t})$ (the second term is regarding as a constant). One can check that $\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{\Theta}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right)$.
Now, let us compute the gradient term. By Eqn. (20) and Eqn. (22), we have
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}}\right]$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r_{t}){\bm{C}}^{\intercal}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}{\bm{C}}^{\intercal}}\right].$
Hence, the update scheme is
${\bm{\Theta}}_{t}=$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$ (24)
$=$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r_{t}){\bm{C}}^{\intercal}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}{\bm{C}}^{\intercal}}\right],$
which is equivalent to
${\bm{W}}_{t}=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$
$=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}}\right],$
which has the same form as Linear-CTD (Eqn. (9)) with the step size $2\alpha\iota_{K}$.
Appendix D Omitted Results and Proofs in Section 4
D.1 Proof of Lemma 5.1
Proof.
By Lemma E.1 and Eqn. (19), we have
$\left({{\mathcal{L}}({\bm{w}})}\right)^{2}=$ $W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}_{{\bm{w}}^{\star}}}\right)$
$\leq$ $\frac{1}{1-\gamma}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}_{{\bm{w}}^{\star}}}\right)$
$=$ $\frac{\iota_{K}}{1-\gamma}\operatorname{tr}\left({\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)}\right)$
$=$ $\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$
We only need to show that
${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq\frac{1}{(1-\sqrt{\gamma})^{2}}\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({\preccurlyeq\frac{4}{(1-\gamma)^{2}\lambda_{\min}}\bar{{\bm{A}}}^{\intercal}\bar{{\bm{A}}}}\right),$ (25)
or equivalently,
$\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}.$
Recall
$\bar{{\bm{A}}}$ $=\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right],$
then for any ${\bm{w}}\in{\mathbb{R}}^{dK}$ with $\left\|{{\bm{w}}}\right\|=1$,
${\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{w}}$
$\qquad=\left\|{\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{w}}}\right\|^{2}$
$\qquad=\left\|{{\bm{w}}-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{w}}}\right\|^{2}$
$\qquad\geq\left({1-\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{w}}}\right\|}\right)^{2}.$
It suffices to show that
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$ (26)
For brevity, we abbreviate ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ as ${\bm{Y}}(r)=\left({y_{ij}(r)}\right)_{i,j=1}^{K}\in{\mathbb{R}}^{K\times K}$. Thus, it suffices to show that
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$
For any vectors ${\bm{w}}=\left({{\bm{w}}(0)^{\intercal},\cdots,{\bm{w}}(K-1)^{\intercal}}\right)$ and ${\bm{v}}=\left({{\bm{v}}(0)^{\intercal},\cdots,{\bm{v}}(K-1)^{\intercal}}\right)$ in ${\mathbb{R}}^{dK}$, we define the corresponding matrices ${\bm{W}}=\left({{\bm{w}}(0),\cdots,{\bm{w}}(K-1)}\right)$ and ${\bm{V}}=\left({{\bm{v}}(0),\cdots,{\bm{v}}(K-1)}\right)$ in ${\mathbb{R}}^{d\times K}$, then $\left\|{{\bm{w}}}\right\|=\left\|{{\bm{W}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{W}}^{\intercal}{\bm{W}}}\right)}$ and $\left\|{{\bm{v}}}\right\|=\left\|{{\bm{V}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{V}}^{\intercal}{\bm{V}}}\right)}$. With these notations, we have
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\bm{w}}^{\intercal}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{v}}$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r){\bm{w}}(i)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)}\right],$
it is easy to check that
${\bm{w}}(i)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)=\left({{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij},$
hence
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r)\left({{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij}}\right]$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\operatorname{tr}\left({{\bm{Y}}(r){\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}}\right)}\right]$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{Y}}(r){\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]$
$\qquad\leq\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{Y}}(r)}\right\|\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,s^{\prime}}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|^{2}}\right]{\mathbb{E}}_{s^{\prime}}\left[{\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|^{2}}\right]}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right]{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\intercal}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\intercal}}\right)}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\left\|{{\bm{W}}}\right\|_{F}\left\|{{\bm{V}}}\right\|_{F}$
$\qquad=\sqrt{\gamma},$
where we used $\left\|{{\bm{Y}}(r)}\right\|\leq\sqrt{\gamma}$ for any $r\in[0,1]$ by Lemma E.3, and Cauchy-Schwarz inequality.
To summarize, we have shown the desired result
$\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-\frac{1}{2}}\bar{{\bm{A}}}^{\intercal}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-1}\bar{{\bm{A}}}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-\frac{1}{2}}\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}.$
D.2 Proof of Lemma 5.2
Proof.
For simplicity, we omit $t$ in the random variables, for example, we use ${\bm{A}}$ to denote a random matrix with the same distribution as ${\bm{A}}_{t}$. In addition, we omit the subscripts in the expectation, the involving random variables are $s\sim\mu_{\pi},a\sim\pi\left({\cdot\mid s}\right),r\sim{\mathcal{P}}_{R}\left({\cdot\mid s,a}\right),s^{\prime}\sim P\left({\cdot\mid s,a}\right)$.
Bounding $C_{A}$.
By Lemma A.3,
$\left\|{{\bm{A}}}\right\|\leq$ $\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right\|+\left\|{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right\|$
$=$ $\left\|{{\bm{\phi}}(s)}\right\|^{2}+\left\|{{\bm{\phi}}(s)}\right\|\left\|{{\bm{\phi}}(s^{\prime})}\right\|\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|$
$\leq$ $1+\sqrt{\gamma},$
where we used Lemma E.3. Hence, $C_{A}\leq 2(1+\sqrt{\gamma})$.
Bounding $C_{e}$.
By Eqn. (29),
$\left\|{{\bm{A}}{\bm{\theta}}^{\star}}\right\|^{2}=$ $\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\bm{A}}^{\intercal}{\bm{A}}{\bm{\theta}}^{\star}$
$\leq$ $2\left({\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}+\gamma\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}}\right)$
$\leq$ $2(1+\gamma)\sup_{s\in{\mathcal{S}}}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}^{2}$
$\leq$ $2(1+\gamma)\sup_{s\in{\mathcal{S}}}\left\|{{\bm{\phi}}(s)}\right\|^{2}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}$
$\leq$ $2(1+\gamma)\left\|{{\bm{\theta}}^{\star}}\right\|^{2}.$
Hence
$\left\|{{\bm{A}}{\bm{\theta}}^{\star}}\right\|\leq$ $\sqrt{2(1+\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|.$
As for $\left\|{{\bm{b}}}\right\|$,
$\left\|{{\bm{b}}}\right\|=$ $\frac{1}{K+1}\left\|{\left[{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s)}\right\|$ (27)
$\leq$ $\frac{1}{K+1}\left\|{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right\|\left\|{{\bm{\phi}}(s)}\right\|$
$\leq$ $\frac{1}{K+1}\left\|{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right\|.$
By Proposition B.3 with ${\bm{\eta}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ satisfying $\eta(\tilde{s})=\nu$ for all $\tilde{s}\in{\mathcal{S}}$, where $\nu=\left({K+1}\right)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ is the discrete uniform distribution, we can derive that, for any $r\in[0,1]$ and $s^{\prime}\in{\mathcal{S}}$, it holds that
$\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)=$ $\left({{\bm{p}}_{\left({b_{r,\gamma}}\right)_{\#}{\eta}(s^{\prime})}-\frac{1}{K+1}\bm{1}_{K}}\right)-\tilde{\bm{G}}(r)\left({{\bm{p}}_{\bm{\eta}}(s^{\prime})-\frac{1}{K+1}\bm{1}_{K}}\right)$
$=$ ${\bm{p}}_{\left({b_{r,\gamma}}\right)_{\#}\nu}-\frac{1}{K+1}\bm{1}_{K},$
Hence,
$\left\|{{\bm{b}}}\right\|\leq$ $\frac{1}{K+1}\left\|{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right\|$
$=$ $\left\|{{\bm{C}}\left({{\bm{p}}_{\left({b_{r,\gamma}}\right)_{\#}\nu}-\frac{1}{K+1}\bm{1}_{K}}\right)}\right\|$
$=$ $\frac{1}{\sqrt{\iota_{K}}}\ell_{2}\left({{\bm{\Pi}}_{K}(b_{r,\gamma})_{\#}(\nu),\nu}\right)$
$\leq$ $\sqrt{K(1-\gamma)}\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$
$\leq$ $3\sqrt{K}(1-\gamma),$
where we used the orthogonal decomposition (Proposition B.2) and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ (Lemma E.4).
In summary,
$\left\|{{\bm{e}}}\right\|=$ $\left\|{{\bm{A}}{\bm{\theta}}^{\star}-{\bm{b}}}\right\|$
$\leq$ $\left\|{{\bm{A}}{\bm{\theta}}^{\star}}\right\|+\left\|{{\bm{b}}}\right\|$
$\leq$ $\sqrt{2(1+\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|+3\sqrt{K}\left({1-\gamma}\right).$
Hence, $C_{e}\leq\sqrt{2(1+\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|+3\sqrt{K}\left({1-\gamma}\right)$.
Bounding $\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)$.
$\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)=$ ${\mathbb{E}}\left[{\left\|{{\bm{A}}{\bm{\theta}}^{\star}-{\bm{b}}}\right\|^{2}}\right]\leq 2\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]{\bm{\theta}}^{\star}+2{\mathbb{E}}\left[{{\bm{b}}^{\intercal}{\bm{b}}}\right].$
By Eqn. (30),
$\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]{\bm{\theta}}^{\star}\leq 2(1+\gamma)\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}.$
And by Lemma E.4,
${\mathbb{E}}\left[{{\bm{b}}^{\intercal}{\bm{b}}}\right]\leq$ $9K(1-\gamma)^{2}.$
To summarize,
$\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)\leq$ $4(1+\gamma)\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}+18K(1-\gamma)^{2}.$
D.3 Proof of Lemma 5.3
Proof.
For simplicity, we use the same abbreviations as in Appendix D.2. As in the proof of [Lemma 2 Samsonov et al., 2024b], we only need to show that, for any $p\in{\mathbb{N}}$, $\alpha\in\left({0,(1-\sqrt{\gamma})/(38p)}\right)$, it holds that
${\mathbb{E}}\left\{{\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\intercal}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)}\right]^{p}}\right\}\preccurlyeq{\bm{I}}_{dK}-\frac{1}{2}\alpha p(1-\gamma){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\left({\preccurlyeq\left({1-\frac{1}{2}\alpha p(1-\gamma)\lambda_{\min}}\right){\bm{I}}_{dK}}\right).$
Let ${\bm{B}}:={\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}$ which satisfies $\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\intercal}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)={\bm{I}}_{dK}-\alpha{\bm{B}}$, we need to bound ${\mathbb{E}}\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{B}}}\right)^{p}}\right]$, it suffices to show that
${\mathbb{E}}\left[{{\bm{B}}}\right]\succcurlyeq(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},\quad{\mathbb{E}}\left[{{\bm{B}}^{p}}\right]\preccurlyeq\frac{17}{16}4^{p}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},\quad\forall p\in\left\{{2,3,\cdots}\right\},$
if we take $\alpha\in\left({0,(1-\sqrt{\gamma})/(2(1+\gamma))}\right)$.
Given these results, we have, when $\alpha\in\left({0,(1-\sqrt{\gamma})/(38p)}\right)$, it holds that
${\mathbb{E}}\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{B}}}\right)^{p}}\right]\preccurlyeq$ ${\bm{I}}-\alpha p{\mathbb{E}}\left[{{\bm{B}}}\right]+\sum_{l=2}^{p}\alpha^{l}\binom{p}{l}{\mathbb{E}}\left[{{\bm{B}}^{l}}\right]$
[…]
$=$ ${\bm{I}}-\left({\alpha p(1-\sqrt{\gamma})-\frac{17\alpha^{2}p^{2}}{1-4\alpha p}}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$
[…]
Lower Bound of ${\mathbb{E}}\left[{{\bm{B}}}\right]$.
To show ${\mathbb{E}}\left[{{\bm{B}}}\right]\succcurlyeq(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, we first show that ${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\succcurlyeq 2(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, which is equivalent to $\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}\succcurlyeq 2(1-\sqrt{\gamma})$, where
${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]$ $=2\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}.$
Then, for any ${\bm{w}}\in{\mathbb{R}}^{Kd}$ with $\left\|{{\bm{w}}}\right\|=1$,
${\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{w}}$
$\qquad=2-2{\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{w}}$
$\qquad\geq 2-2\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad\geq 2(1-\sqrt{\gamma}),$
where we used the result Eqn. (26).
Next, we give an upper bound for ${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]$, we need to compute the following terms: by Lemma A.4,
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}=$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$ (28)
$=$ $\left\|{{\bm{\phi}}(s)}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
[…]
Hence
${\mathbb{E}}\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$
And by Lemma A.4 and Lemma E.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]$
$\qquad=\left({{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s)}\right\|^{2}\left({{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right).$
To summarize, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\intercal}\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\preccurlyeq 2\left({{\bm{B}}_{1}^{\intercal}{\bm{B}}_{1}+{\bm{B}}_{2}^{\intercal}{\bm{B}}_{2}}\right)$, we have
${\bm{A}}^{\intercal}{\bm{A}}\preccurlyeq$ $2{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right),$ (29)
and, after taking expectation,
${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (30)
Combine these together, we obtain
${\mathbb{E}}\left[{{\bm{B}}}\right]$ $\succcurlyeq 2\left[{(1-\sqrt{\gamma})-\alpha(1+\gamma)}\right]{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\succcurlyeq(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},$
if we take $\alpha\in\left({0,(1-\sqrt{\gamma})/(2(1+\gamma))}\right)$.
Upper Bound of ${\mathbb{E}}\left[{{\bm{B}}^{p}}\right]$.
Because ${\bm{B}}^{2}$ is always PSD, we have the following upper bound
${\bm{B}}^{p}\preccurlyeq\left\|{{\bm{B}}}\right\|^{p-2}{\bm{B}}^{2}.$
We first give an almost-sure upper bound for $\left\|{{\bm{B}}}\right\|$. By Lemma 5.2, $\left\|{{\bm{A}}}\right\|\leq 1+\sqrt{\gamma}$. And by Eqn. (29),
$\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|\leq$ $2\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right\|$ (31)
$\leq$ $2\left\|{{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right\|$
$\leq$ $2(1+\gamma).$
Hence,
$\left\|{{\bm{B}}}\right\|=$ $\left\|{{\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}}\right\|$ (32)
$\leq$ $2\left\|{{\bm{A}}}\right\|+\alpha\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|$
$\leq$ $2(1+\sqrt{\gamma})+2\alpha(1+\gamma)$
$\leq$ $4,$
because $\alpha\in\left({0,(1-\sqrt{\gamma})/(2(1+\gamma))}\right)$.
Now, we aim to give an upper bound for ${\mathbb{E}}\left[{{\bm{B}}^{2}}\right]$,
${\bm{B}}^{2}=$ $\left({{\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}$
[…]
where we used the fact that $\left({{\bm{B}}_{1}+{\bm{B}}_{2}}\right)^{2}\preccurlyeq(1+\beta){\bm{B}}_{1}^{2}+(1+\beta^{-1}){\bm{B}}_{2}^{2}$ for any symmetric matrices ${\bm{B}}_{1},{\bm{B}}_{2}$, since $\beta{\bm{B}}_{1}^{2}+\beta^{-1}{\bm{B}}_{2}^{2}-{\bm{B}}_{1}{\bm{B}}_{2}-{\bm{B}}_{2}{\bm{B}}_{1}=\left({\sqrt{\beta}{\bm{B}}_{1}-\sqrt{\beta^{-1}}{\bm{B}}_{2}}\right)^{2}\succcurlyeq\bm{0}$, $\beta\in(0,1)$ to be determined; and the fact that ${\bm{A}}^{2}+\left({{\bm{A}}^{\intercal}}\right)^{2}\preccurlyeq{\bm{A}}^{\intercal}{\bm{A}}+{\bm{A}}{\bm{A}}^{\intercal}$ since the square of the skew-symmetric matrix is negative semi-definite $\left({{\bm{A}}-{\bm{A}}^{\intercal}}\right)^{2}\preccurlyeq\bm{0}$. By Eqn. (31) and Eqn. (30), we have
$\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|\leq 2(1+\gamma),$
${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},$ (33)
thus, by $\alpha\in\left({0,(1-\sqrt{\gamma})/(2(1+\gamma))}\right)$, it holds that
$\alpha^{2}{\mathbb{E}}\left[{\left({{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}}\right]\preccurlyeq 4\alpha^{2}(1+\gamma)^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq(1-\sqrt{\gamma})^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (34)
As for ${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\intercal}}\right]$, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\intercal}\preccurlyeq 2\left({{\bm{B}}_{1}{\bm{B}}_{1}^{\intercal}+{\bm{B}}_{2}{\bm{B}}_{2}^{\intercal}}\right)$, we have
${\bm{A}}{\bm{A}}^{\intercal}=$$\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]}\right\}$
$\cdot\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]}\right\}^{\intercal}$
[…]
By Eqn. (28), we have
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right).$
And by Lemma E.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}$
$\qquad=\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s^{\prime})}\right\|^{2}\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right).$
To summarize, we have
${\bm{A}}{\bm{A}}^{\intercal}\preccurlyeq$ $2(1+\gamma){\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right),$
after taking expectation
${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\intercal}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (35)
Put everything together (Eqn. (33), Eqn. (34), Eqn. (35)), we have
${\bm{B}}^{2}\preccurlyeq$ $2(1+\beta)\left({{\bm{A}}^{\intercal}{\bm{A}}+{\bm{A}}{\bm{A}}^{\intercal}}\right)+(1+\beta^{-1})\alpha^{2}\left({{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}$
[…]
where we take $\beta=\frac{1-\sqrt{\gamma}}{\sqrt{8(1+\gamma)}}$. Therefore, by Eqn. (32),
${\bm{B}}^{p}\preccurlyeq$ $\left\|{{\bm{B}}}\right\|^{p-2}{\bm{B}}^{2}$
[…]
$=$ $\frac{17}{16}4^{p}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$
D.4 Proof of Theorem 4.1
Proof.
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with [Theorem 1 Samsonov et al., 2024b], we have
${\mathbb{E}}^{1/2}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{2}}\right]\lesssim$ $\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\sqrt{\frac{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}{T}}\left({1+\frac{C_{A}\sqrt{\alpha}}{\sqrt{a}}}\right)+\frac{\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}{\sqrt{\alpha a}T}$
$+(1-\alpha a)^{T/2}\left({\frac{1}{\alpha T}+\frac{C_{A}}{\sqrt{\alpha a}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\lesssim$ $\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{T}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)$
$+\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{\alpha(1-\gamma)\lambda_{\min}}T}$
$+(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}\left({\frac{1}{\alpha T}+\frac{1}{\sqrt{\alpha(1-\gamma)\lambda_{\min}}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\lesssim$ $\frac{1}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)$
$+\frac{1}{T}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{1}{T}\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}+\frac{1}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|.$
D.5 $L^{p}$ Convergence
Theorem D.1 ($L^{p}$ Convergence).
For any $K\geq(1-\gamma)^{-1}$, $p>2$, $T\geq 2$, $\alpha\in(0,(1-\sqrt{\gamma})/[38(p+\log T)])$ and initialization ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$, it holds that
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{p}}\right]\lesssim$ $\frac{\sqrt{p}}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\frac{\sqrt{\alpha p}+\alpha p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)$
$+\frac{p}{T}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}\left({1+\frac{1}{\sqrt{\alpha p}}}\right)$
$+\frac{1}{T}\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}+\frac{p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|.$
Proof.
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with [Theorem 2 Samsonov et al., 2024b], we have
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{p}}\right]$
$\qquad\lesssim\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\sqrt{\frac{p\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}{T}}\left({1+\frac{C_{A}\sqrt{\alpha p}}{\sqrt{a}}+\frac{C_{A}C_{e}\alpha p}{\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}}\right)+\frac{(1+C_{A})C_{e}p}{T}$
$\qquad\quad+\frac{p\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}{\sqrt{a}T}\left({1+\frac{1}{\sqrt{\alpha p}}}\right)+(1-\alpha a)^{T/2}\left({\frac{1}{\alpha T}+\frac{C_{A}p}{\sqrt{\alpha a}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\qquad\lesssim\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\sqrt{p}\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{T}}\left({1+\frac{\sqrt{\alpha p}+\alpha p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)+\frac{p\left({\left\|{{\bm{\theta}}^{\star}}\right\|+\sqrt{K}\left({1-\gamma}\right)}\right)}{T}$
$\qquad\quad+p\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{(1-\gamma)\lambda_{\min}}T}\left({1+\frac{1}{\sqrt{\alpha p}}}\right)$
$\qquad\quad+(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}\left({\frac{1}{\alpha T}+\frac{p}{\sqrt{\alpha(1-\gamma)\lambda_{\min}}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\qquad\lesssim\frac{\sqrt{p}}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\frac{\sqrt{\alpha p}+\alpha p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)$
$\qquad\quad+\frac{p}{T}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}\left({1+\frac{1}{\sqrt{\alpha p}}}\right)$
$\qquad\quad+\frac{1}{T}\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}+\frac{p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|,$
where we used the fact that $\left\|{{\bm{\theta}}^{\star}}\right\|\leq\left({\lambda_{\min}}\right)^{-1/2}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$. ∎
D.6 Convergence Results for SSGD with the PMF Representation
In this section, we present without proof, the counterparts of Lemma 5.1, Lemma 5.2 and Lemma 5.3 for SSGD with the PMF representation. These results will additionally depend on $K$. The additional $K$-dependent terms arise because the condition number of ${\bm{C}}^{\intercal}{\bm{C}}$ scales with $K^{2}$ (Lemma E.2). These terms are inevitable within our theoretical framework. By [Theorem 1 Samsonov et al., 2024b], we find that we cannot obtain a convergence rate independent of $K$ for the vanilla form Eqn. (23) without a preconditioner as in our Linear-CTD. A natural question is whether the convergence rate of this algorithm truly depends on $K$. We leave this as future work.
The proofs of these results only require minor modifications to the original proofs. For the sake of simplicity, we will not introduce new notations. Instead, we will only add the subscript $\operatorname{PMF}$ to the original notations to indicate the difference.
According to Eqn. (23) in Appendix 7, the algorithm corresponds to the following linear system for ${\bm{w}}\in{\mathbb{R}}^{dK}$
$\bar{{\bm{A}}}_{\operatorname{PMF}}{\bm{w}}=\bar{{\bm{b}}}_{\operatorname{PMF}},$
where
$\bar{{\bm{A}}}_{\operatorname{PMF}}$ $=\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r)}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right],$
$\bar{{\bm{b}}}_{\operatorname{PMF}}=\frac{1}{K+1}{\mathbb{E}}_{s,r}\left\{{\left[{{\bm{C}}^{\intercal}{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s)}\right\}.$
Lemma D.1.
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$, it holds that
$\left({{\mathcal{L}}({\bm{w}})}\right)^{2}\leq\frac{16}{K(1-\gamma)^{4}\lambda_{\min}}\left\|{\bar{{\bm{A}}}_{\operatorname{PMF}}\left({{{\bm{w}}}-{\bm{w}}^{\star}}\right)}\right\|^{2}.$
Lemma D.2.
$C_{A,\operatorname{PMF}}\leq 4K^{2},$
$C_{e,\operatorname{PMF}}\leq 4K^{\frac{3}{2}}\left({\sqrt{K}\left\|{{\bm{w}}^{\star}}\right\|+\left({1-\gamma}\right)}\right),$
$\operatorname{tr}\left({{\bm{\Sigma}}_{e,\operatorname{PMF}}}\right)\leq 18K^{2}\left({\left\|{{\bm{w}}^{\star}}\right\|_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}+K(1-\gamma)^{2}}\right).$
Lemma D.3.
For any $p\geq 2$, let $a=(1-\sqrt{\gamma})\lambda_{\min}/2$ and $\alpha_{p,\infty}=(1-\sqrt{\gamma})/(38K^{2})$ ($\alpha_{p,\infty}p\leq 1/2$). Then for any $\alpha\in\left({0,\alpha_{p,\infty}}\right)$, ${\bm{u}}\in{\mathbb{R}}^{dK}$ and $n\in{\mathbb{N}}$
${\mathbb{E}}^{1/p}\left[{\left\|{{\bm{\Gamma}}_{1:n,\operatorname{PMF}}^{(\alpha)}{\bm{u}}}\right\|^{p}}\right]\leq\left({1-\alpha a}\right)^{n}\left\|{{\bm{u}}}\right\|.$
Appendix E Other Technical Lemmas
Lemma E.1.
For any $\nu_{1},\nu_{2}\in{\mathscr{P}}^{\mathsf{sign}}$, we have $W_{1}(\nu_{1},\nu_{2})\leq\frac{1}{\sqrt{1-\gamma}}\ell_{2}(\nu_{1},\nu_{2})$.
Proof.
By Cauchy-Schwarz inequality,
[…]
$=$ $\frac{1}{\sqrt{1-\gamma}}\ell_{2}(\nu_{1},\nu_{2}).$
Lemma E.2.
Let ${\bm{C}}\in{\mathbb{R}}^{K\times K}$ be the matrix defined in Eqn. (16), it holds that the eigenvalues of ${\bm{C}}^{T}{\bm{C}}$ are $1/(4\cos^{2}(k\pi/(2K+1))$ for $k\in[K]$ , and thus
$\left\|{{\bm{C}}^{\intercal}{\bm{C}}}\right\|=\frac{1}{4\sin^{2}\frac{\pi}{4K+2}}\leq K^{2},\quad\left\|{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1}}\right\|=4\sin^{2}\frac{K\pi}{4K+2}\leq 4.$ (36)
Proof.
One can check that
${\bm{C}}^{\intercal}{\bm{C}}=\begin{bmatrix}K&K-1&\cdots&2&1\\ K-1&K-1&\cdots&2&1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 2&2&\cdots&2&1\\ 1&1&\cdots&1&1\end{bmatrix},$
$\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1}=\begin{bmatrix}1&-1&0&\cdots&0&0\\ -1&2&-1&\cdots&0&0\\ 0&-1&2&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&2&-1\\ 0&0&0&\cdots&-1&1\end{bmatrix}.$
Then, one can work with the the inverse of ${\bm{C}}^{\intercal}{\bm{C}}$ and calculate its singular values by induction, which has similar forms to the analysis of Toeplitz’s matrix. See [Godsil, 1985] for more details. ∎
Lemma E.3.
For any $r\in[0,1]$, it holds that $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|\leq\sqrt{\gamma}$ and $\left\|{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{1/2}\tilde{{\bm{G}}}(r)\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1/2}}\right\|\leq\sqrt{\gamma}$.
Proof.
One can check that
${\bm{C}}^{-1}=\begin{bmatrix}1&0&\cdots&0&0\\ -1&1&\cdots&0&0\\ 0&-1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-1&1\end{bmatrix}.$
It is clear that
$\left\|{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{1/2}\tilde{{\bm{G}}}(r)\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1/2}}\right\|\leq\sqrt{\gamma}$ $\iff{\bm{Y}}(r){\bm{Y}}^{\intercal}(r)\preccurlyeq\gamma{\bm{I}}$ (37)
$\iff\tilde{{\bm{G}}}(r)({\bm{C}}^{\intercal}{\bm{C}})^{-1}\tilde{{\bm{G}}}^{\intercal}(r)\preccurlyeq\gamma({\bm{C}}^{\intercal}{\bm{C}})^{-1}$
$\iff{\bm{C}}\tilde{{\bm{G}}}(r)({\bm{C}}^{\intercal}{\bm{C}})^{-1}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}\preccurlyeq\gamma{\bm{I}}$
$\iff\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|\leq\sqrt{\gamma}.$
By Lemma F.2 and an upper bound on the spectral norm (Riesz–Thorin interpolation theorem) [Theorem 7.3 Serre, 2002], we obtain that
$\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|\leq\sqrt{\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{1}\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{\infty}}\leq\sqrt{1\cdot\gamma}=\sqrt{\gamma}.$ (38)
Lemma E.4.
Suppose $K\geq(1-\gamma)^{-1}$, $\nu=(K+1)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ is the discrete uniform distribution, then for any $r\in[0,1]$, it holds that
$\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)\leq 3\sqrt{1-\gamma}.$
Proof.
Let $\tilde{\nu}$ be the continuous uniform distribution on $\left[{0,(1-\gamma)^{-1}+\iota_{K}}\right]$, we consider the following decomposition
$\ell_{2}\left({\nu,(b_{r,\gamma})_{\#}(\nu)}\right)\leq\ell_{2}\left({\nu,\tilde{\nu}}\right)+\ell_{2}\left({\tilde{\nu},(b_{r,\gamma})_{\#}(\tilde{\nu})}\right)+\ell_{2}\left({(b_{r,\gamma})_{\#}(\tilde{\nu}),(b_{r,\gamma})_{\#}(\nu)}\right).$
By definition, we have
[…]
$=$ $\sqrt{\frac{1}{3K(K+1)(1-\gamma)}}$
$\leq$ $\frac{1}{K\sqrt{1-\gamma}}.$
By the contraction property, we have
$\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),(b_{r,\gamma})_{\#}(\tilde{\nu})}\right)\leq$ $\sqrt{\gamma}\ell_{2}\left({\nu,\tilde{\nu}}\right)\leq\frac{\sqrt{\gamma}}{K\sqrt{1-\gamma}}.$
We only need to bound $\ell_{2}\left({\tilde{\nu},(b_{r,\gamma})_{\#}(\tilde{\nu})}\right)$. We can find that $(b_{r,\gamma})_{\#}(\tilde{\nu})$ is the continuous uniform distribution on $\left[{r,r+\gamma\iota_{K}+\gamma(1-\gamma)^{-1}}\right]$, and the upper bound is less than the upper bound of $\nu$, namely, $r+\gamma\iota_{K}+\gamma(1-\gamma)^{-1}\leq(1-\gamma)^{-1}+\gamma\iota_{K}<(1-\gamma)^{-1}+\iota_{K}$. Hence
[…]
$=$ $\frac{(1-\gamma)^{2}K^{2}r^{3}}{3(K+1)^{2}}+\left({\frac{(1-\gamma)\gamma K^{2}r^{3}}{3(K+1)^{2}}+\frac{(1-\gamma)\gamma K^{2}\left({\frac{K+1}{K}-r}\right)^{3}}{3(K+1)^{2}}}\right)+\frac{(1-\gamma)^{2}K^{2}\left({\frac{K+1}{K}-r}\right)^{3}}{3(K+1)^{2}}$
$\leq$ $(1-\gamma)^{2}+(1-\gamma)\gamma$
$=$ $1-\gamma.$
To summarize, we have
$\ell_{2}\left({\nu,(b_{r,\gamma})_{\#}(\nu)}\right)\leq\frac{1}{K\sqrt{1-\gamma}}+\sqrt{1-\gamma}+\frac{\sqrt{\gamma}}{K\sqrt{1-\gamma}}\leq 3\sqrt{1-\gamma},$
where we used the assumption $K\geq(1-\gamma)^{-1}$. ∎
Appendix F Analysis of the Categorical Projected Bellman Matrix
Recall that $\tilde{{\bm{G}}}(r)={\bm{G}}(r)-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r)$. We extend the definition in Theorem 3.1 and let $g_{j,k}(r)=h\left({(r+\gamma x_{j}-x_{k})/\iota_{K}}\right)_{+}=h(r/\iota_{K}+\gamma j-k)$ for $j,k\in\{0,1,\cdots,K\}$ where $h(x)=\left({1-\left|{x}\right|}\right)_{+}$.
Lemma F.1.
For any $r\in[0,1]$ and any $k\in\{0,1,\cdots,K\}$, in ${\bm{g}}_{k}(r)$ there is either only one nonzero entry or two adjacent nonzero entries.
Proof.
It is clear that $h(x)>0\iff-1<x<1$. Let $k_{j}(r)=\min\ \{k:g_{j,k}(r)>0\}$, then $k_{j}(r)=\min\{k:r/\iota_{K}+\gamma j-k<1\}=\min\{k:0\leq r/\iota_{K}+\gamma j-k<1\}$. The existence of $k_{j}(r)$ is due to
$r/\iota_{K}+\gamma j-K\leq 1/\iota_{K}+\gamma j-K\leq(1-\gamma)K+\gamma K-K=0<1.$ (39)
Let $a_{j}(r):=r/\iota_{K}+\gamma j-k_{j}(r)\in[0,1)$. Then $g_{j,k_{j}(r)}(r)=h(a_{j}(r))=1-a_{j}(r)$ and $g_{j,k_{j}(r)+1}(r)=h(a_{j}(r)-1)=a_{j}(r)$ are the only entries that can be nonzeros. ∎
The following results are immediate corollaries.
Corollary F.1.
$\sum_{k=0}^{i}g_{j,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 1,&\text{ for}\ k_{j}(r)<i\leq K.\\ \end{cases}$ (40)
Corollary F.2.
For any $\gamma<1$, it holds that
$k_{j+1}(r)=\begin{cases}k_{j}(r),&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ k_{j}(r)+1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$ (41)
As a result,
$a_{j+1}(r)=\begin{cases}a_{j}(r)+\gamma,&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ a_{j}(r)+\gamma-1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$ (42)
Lemma F.2.
All entries in ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ are non-negative. $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{\infty}=\gamma$ and $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{1}\leq 1$.
Proof.
By definition the entries of $\tilde{{\bm{G}}}(r)$ are
$(\tilde{{\bm{G}}}(r))_{j,i}=g_{j,i}(r)-g_{K,i}(r)\quad\text{for }j,i\in\{0,1,\cdots,K-1\}.$ (43)
Using the previous corollaries, direct calculation show that if $k_{j+1}(r)=k_{j}(r)$,
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ a_{j+1}(r)-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$ (44)
And if $k_{j+1}(r)=k_{j}(r)+1$,
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0<i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ a_{j+1}(r),&\text{ for}\ i=k_{j+1}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$ (45)
As a result, all entries in ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is non-negative. Moreover, the sum of each column and $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{\infty}$ is $\gamma$ because
$\sum_{i=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\begin{cases}a_{j+1}(r)-a_{j}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r),\\ 1-a_{j}(r)+a_{j+1}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r)+1.\\ \end{cases}$ (46)
The row sum of ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is
$\sum_{j=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{m=0}^{i}g_{0,m}(r)-\sum_{m=0}^{i}g_{K,m}(r)\leq 1-0=1.$ (47)
Thus, it holds that $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{1}\leq 1$. ∎